### v2

# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

Abstract
Fully decentralized training of machine learning models offers significant advantages in scalability, robustness, and fault tolerance. However, achieving differential privacy (DP) guarantees in such settings is challenging due to the absence of a central aggregator and varying trust assumptions among nodes. We present a novel privacy analysis of decentralized gossip-based averaging algorithms with additive node-level noise, from arbitrary views of nodes in a graph and especially consider the averaging over nearest neighbors with secure summation and individual node-wise views. Our main contribution is a new analytical framework based on a linear systems formulation that accurately characterizes privacy leakage between nodes across different scenarios. In case the gossip averaging happens via secure summation, we show that the Rényi DP parameter growth is asymptotically $O(T)$, where $T$ is the number of training rounds, similarly as in the case of central aggregation.
1 Introduction
Common federated learning (FL) scenarios assume the presence of a central parameter server for coordinating model updates. In contrast, fully decentralized setups operate without a central orchestrator: compute nodes, each holding a private dataset, directly exchange model states or updates with a subset of peers. Such decentralized architectures offer advantages in scalability, fault tolerance, and robustness, but also introduce new algorithmic and privacy challenges.
Fully decentralized gradient-based optimization methods are typically distributed variants of gradient descent and can be broadly classified into two categories.
In random walk-based methods (rw1; rw2; rw3), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix. The neighbor then updates its local parameters and passes the updated model to another randomly chosen neighbor. Over time, and under mild assumptions on the transition matrix and the network graph (e.g., connectedness, symmetry), the random walk ensures uniform coverage of all nodes. However, such sequential update dynamics inherently limits parallelism and scalability.
In gossip averaging-based methods (gossip2; gossip1), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states. Each node updates its parameters by averaging over the received messages. These operations are repeated iteratively until the network reaches approximate consensus on the model parameters. Gossip protocols are naturally parallelizable and more scalable than random walk approaches, making them attractive for large-scale decentralized learning. This work focuses on such gossip-based protocols for DP decentralized learning.
DP in Centralized FL. In centralized FL (FLbook), distributed DP (UllahCKO23; google_flblog_2023; FLDP3) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (truex2019hybrid; Erlingsson2019; bell2020secure). This not only eliminates the need to trust individual nodes but also enables lower per-user noise by aggregating over multiple contributions.
DP in Decentralized Optimization. The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (dlattack4; dlattack3; dlattack1; dlattack2). Early decentralized DP methods achieve privacy by locally perturbing gradients or models (HMV2015; BGTT18; XZW2022), relying on local DP guarantees. This often imposes a poor utility-privacy tradeoff compared to centralized DP (ChanSS12).
To address this, recent works propose pair-wise network DP (networkdp3; networkdp2; networkdp1), where privacy loss is analyzed between pairs of nodes. These relaxations give better utility while accounting for the structure of decentralized communication. networkdp2 give Rényi differential privacy (RDP) bounds that take into account the limited view, however they exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of compositions (or training iterations in ML model training), making them unsuitable for practical ML model training with reasonable privacy guarantees. Recently, bellet2025unified considered the linear-systems viewpoint of decentralized optimization to cover adaptive compositions through the so called matrix mechanisms (kairouz2021practical; pillutla2025correlated) which leads to considerably more accurate privacy analysis than that of networkdp2. Nevertheless, as our asymptotic sensitivity result (our Theorem 13) shows, the full linear-systems framework is still a valuable tool for understanding the privacy-preserving properties of decentralized optimization.
Our Approach. In this work, we develop techniques to analyze network DP guarantees under various threat models, showing that the total sensitivity grows asymptotically as $O(\sqrt{T})$, resulting in asymptotic $O(T)$ growth of the RDP parameters, for example, representing a significant improvement over prior analyses. This network DP accounting is obtained by interpreting the dynamics of the gossip averaging as a linear state-space system and the view of individual nodes or subsets of nodes as projected Gaussian mechanisms (GMs). This also provides tools to analyze the network DP guarantees in different threat scenarios. In addition to considering the threat model of (networkdp2), where the neighboring nodes see the plain messages sent by their neighbors, we also analyze algorithms that incorporate a secure summation protocol between neighboring nodes.
Secure Aggregation in Decentralized Learning. While the amplification effect of summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored. In current gossip-based systems (networkdp2), each node receives messages directly from its neighbors and can, in principle, view their full content. As we show, summation over a node’s neighborhood can give a meaningful privacy amplification. Several studies have proposed integrating MPC primitives into decentralized learning workflows (jayaraman2018distributed; lian2018asynchronous; jeon2021privacy). Concrete protocols for decentralized secure aggregation have been developed in recent works (sabater2022gopa; dlsa1; biswas2024). The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (nicolas2024differentially).
Our Contributions:
  • We provide an analytical framework for evaluating differential privacy guarantees in decentralized gossip-based averaging algorithms. Our framework leverages a linear dynamical systems perspective to track sensitivity propagation through iterative updates.
  • We demonstrate that the DP guarantees can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed with the help of our linear system model.
  • We show that in case secure summation is used to carry averaging over the nearest neighbors, the squared sensitivity scales asymptotically as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on RDP (networkdp2).
2 Background
We first shortly review the required technicalities on differential privacy. We then discuss the projected Gaussian mechanism central to our analysis and define the model for decentralized learning and the network DP guarantees.
2.1 Differential Privacy
An input dataset containing $n$ data points is denoted as $D=(x_{1},\ldots,x_{n})\in\mathcal{D}$, where $\mathcal{D}$ denotes the set of datasets of all sizes. We say that two datasets $D$ and $D^{\prime}$ are neighbors if we get one by adding or removing one element to/from the other (denoted $D\sim D^{\prime}$). We say that a mechanism $\mathcal{M}\,:\,\mathcal{D}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if the output distributions for neighboring datasets are always $(\varepsilon,\delta)$-indistinguishable.
Definition 1 (dwork_et_al_2006).
Let $\varepsilon\geq 0$ and $\delta\in[0,1]$. A randomized mechanism $\mathcal{M}\,:\,\mathcal{D}\rightarrow\mathcal{P}(\mathcal{O})$, where $\mathcal{P}(\mathcal{O})$ denotes the set of probability distributions over the output space $\mathcal{O}$, is $(\varepsilon,\delta)$-DP if for every pair of neighboring datasets $D,D^{\prime}$, every measurable set $O\subset\mathcal{O}$,
$\mathbb{P}(\mathcal{M}(D)\in O)\leq{\rm e}\hskip 1.0pt^{\epsilon}\mathbb{P}(\mathcal{M}(D^{\prime})\in O)+\delta.$
Given two distributions $(P,Q)$, if the conditions $\mathbb{P}(P\in O)\leq{\rm e}\hskip 1.0pt^{\epsilon}\mathbb{P}(Q\in O)+\delta$ and $\mathbb{P}(Q\in O)\leq{\rm e}\hskip 1.0pt^{\epsilon}\mathbb{P}(P\in O)+\delta$ hold for every measurable set $O\subset\mathcal{O}$, we also denote $P\simeq_{(\varepsilon,\delta)}Q$.
With the hockey-stick divergence we can equivalently measure the $(\varepsilon,\delta)$-distance of distributions. For probability distributions $P$ and $Q$, and for $\varepsilon\in\mathbb{R}$, it is defined as $H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(P||Q)=\int[P(t)-{\rm e}\hskip 1.0pt^{\varepsilon}Q(t)]_{+}$, where $z]_{+}=\max\{0,z\}$. The $(\varepsilon,\delta)$-DP guarantees can be then given as follows.
Lemma 2 (balle2018subsampling).
A mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP if and only if, $\max_{D\sim D^{\prime}}H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(\mathcal{M}(X)||\mathcal{M}(X^{\prime}))\leq\delta$.
The Gaussian Mechanism is a common way to achieve $(\epsilon,\delta)$-differential privacy by adding Gaussian noise to a functions output.
Definition 3 (Gaussian Mechanism).
Let $f:\mathcal{D}\to\mathbb{R}^{d}$ be a function with $\ell_{2}$-sensitivity defined as $\Delta_{2}(f)=\max_{D,D^{\prime}}\|f(D)-f(D^{\prime})\|_{2}$, where the maximum is over all adjacent datasets $D,D^{\prime}$ (i.e., datasets differing in at most one individual’s data). The Gaussian mechanism outputs
$\mathcal{M}(D)=f(D)+Z,$
where $Z\sim\mathcal{N}(0,\sigma^{2}I_{d})$.
Running a DP training algorithm for $T$ iterations is commonly modeled as an adaptive composition of $T$ mechanisms such that the adversary has a view on the output of all intermediate outputs. This means that we then analyze mechanisms of the form
$\mathcal{M}^{(T)}(D)=\big(\mathcal{M}_{1}(D),\mathcal{M}_{2}(\mathcal{M}_{1}(D),D),\ldots,\mathcal{M}_{T}(\mathcal{M}_{1}(D),\ldots,\mathcal{M}_{T-1}(D),D)\big).$
The results of (balle2018gauss) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism. Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by dong2022gaussian.
Lemma 4 (dong2022gaussian).
Consider an adaptive composition of $T$ Gaussian mechanisms, each with $L_{2}$-sensitivity $\Delta$ and noise scale parameter $\sigma$. The adaptive composition is $(\varepsilon,\delta)$-DP for
$\delta(\varepsilon)=\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}\cdot\Delta}+\frac{\sqrt{T}\cdot\Delta}{2\sigma}\right)-e^{\varepsilon}\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}\cdot\Delta}-\frac{\sqrt{T}\cdot\Delta}{2\sigma}\right).$
2.2 Projected Gaussian Mechanism and Moore–Penrose Pseudoinverse
In our results, the network DP guarantees become those of a projected Gaussian mechanism of the form
$\mathcal{M}(D)=f(D)+AZ,$ (2.1)
where $f:\mathcal{D}\rightarrow\mathbb{R}^{m}$, $A\in\mathbb{R}^{m\times n}$ and $Z\sim\mathcal{N}(0,\sigma^{2}I_{n})$ for some $\sigma>0$. To analyze the DP guarantees of these mechanism, we will use the Moore–Penrose pseudoinverse of $A$.
We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (golub2013matrix).
Definition 5 (Compact SVD and Moore–Penrose Pseudoinverse).
Let $A\in\mathbb{R}^{m\times n}$ be a matrix of rank $r$. The compact SVD (also known as the economy SVD) of $A$ is given by:
$A=U_{r}\Sigma_{r}V_{r}^{\top}$
where $U_{r}\in\mathbb{R}^{m\times r}$ and $V_{r}\in\mathbb{R}^{n\times r}$ contain the left and right singular vectors corresponding to the non-zero singular values, respectively, and $\Sigma_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the non-zero singular values $\sigma_{1}\geq\dots\geq\sigma_{r}>0$. The Moore–Penrose pseudoinverse of $A$, denoted $A^{+}$, is defined via compact SVD as
$A^{+}=V_{r}\Sigma_{r}^{+}U_{r}^{\top}$
where $\Sigma_{r}^{+}\in\mathbb{R}^{r\times r}$ is a diagonal matrix with entries $\frac{1}{\sigma_{1}},\dots,\frac{1}{\sigma_{r}}$, i.e., $\Sigma_{r}^{+}=\mathrm{diag}\left(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}},\dots,\frac{1}{\sigma_{r}}\right).$
To analyze the gossip averaging algorithms, the following technical result will play a central role.
Lemma 6.
Let $\sigma>0$ and $A\in\mathbb{R}^{m\times n}$. Suppose $\mathcal{M}(D)$ is a projected Gaussian mechanism of the form given in Eq. (2.1). Let $D$ and $D^{\prime}$ be two datasets such that $f(D)-f(D^{\prime})\in\mathrm{Range(A)}$, where $\mathrm{Range(A)}$ denotes the subspace spanned by the columns of $A$. Then, for all $\alpha\geq 0$,
$H_{\alpha}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)=H_{\alpha}\left(\mathcal{N}\big(\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2},\sigma^{2}\big)\,||\,\mathcal{N}(0,\sigma^{2})\right).$
i.e., the $(\varepsilon,\delta)$-distance between $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$ is upper bounded by the $(\varepsilon,\delta)$-DP guarantee of a gaussian mechanism with $L_{2}$-sensitivity $\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2}$ and noise parameter $\sigma$.
2.3 Decentralized Learning and Network Differential Privacy
We consider a connected undirected graph $G=(V,E)$ with $|V|=n$ nodes. The graph structure is encoded by an adjacency matrix
$A\in\{0,1\}^{n\times n},\qquad A_{ij}=1\;\text{iff}\;(i,j)\in E.$
We denote the nodes by indices $n]=\{1,\dots,n\}$. For each node $i\in[n]$, the neighborhood and closed neighborhood are defined as
$N_{i}=\{j\in[n]:(i,j)\in E\},\qquad\bar{N}_{i}=N_{i}\cup\{i\}.$
Local Data and Learning Objective.
The global dataset is partitioned across nodes:
$D=\bigcup_{i=1}^{n}D_{i},$
where node $i$ holds the shard $D_{i}$. The goal of decentralized learning is to minimize the global loss
$f(\theta)=\sum_{i=1}^{n}f_{i}(\theta),\qquad\theta\in\mathbb{R}^{d},$
where $f_{i}(\theta)$ depends only on $D_{i}$. At each round $t\in[T]$, node $i$ maintains a state variable $\theta_{t}]_{i}$. For the analysis, we consider the scalar case ($d=1$); the multidimensional case follows from the matrix mechanisms analogy given by bellet2025unified as described in Section 3.4.
Dataset Adjacency.
We adopt a record-level adjacency relation. Two datasets $D=\cup_{i}D_{i}$ and $D^{\prime}=\cup_{i}D_{i}^{\prime}$ are adjacent, written $D\sim D^{\prime}$, if they differ in exactly one node’s data:
$D\sim_{i}D^{\prime}\quad\iff\quad D_{j}=D^{\prime}_{j}\ \text{for all }j\neq i.$
Mechanism and Observations.
A decentralized mechanism is a randomized mapping
$\mathcal{M}:D\mapsto\text{(all messages exchanged during the protocol)}.$
For a subset of nodes $\mathcal{A}\subseteq[n]$, the view of $\mathcal{A}$ is denoted
$\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A}).$
To extract node-wise components, we define the selector matrix: for $\mathcal{A}=\{i_{1},\dots,i_{|\mathcal{A}|}\}$,
$S(\mathcal{A})=\sum_{k=1}^{|\mathcal{A}|}e_{k}e_{i_{k}}^{\top}\in\mathbb{R}^{|\mathcal{A}|\times n},$
where $e_{k}$ is the $k$th standard basis vector.
Network Differential Privacy.
Node $j$ is $(\varepsilon,\delta)$-DP from the point of view of $\mathcal{A}\subseteq[n]$ if for all adjacent datasets $D\sim_{j}D^{\prime}$,
$\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})\;\simeq_{(\varepsilon,\delta)}\;\mathrm{View}_{\mathcal{M}(D^{\prime})}(\mathcal{A}).$
We assume honest-but-curious nodes: they follow the protocol but analyze observed messages. As in (networkdp2), we first allow noise added by observers to contribute to privacy; this assumption is removed in Section 3.3.
Gossip-Based Averaging.
Let $x_{t}\in\mathbb{R}^{n}$ denote the vector of local values exchanged at time $t$. A gossip update is given by
$\theta_{t+1}=Wx_{t},$
where $W\in\mathbb{R}^{n\times n}$ is the gossip matrix. We call $W$:
$\text{row-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1},\qquad\text{doubly-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1}\ \text{and}\ W^{\top}\mathbf{1}=\mathbf{1}.$
For the error analysis we also consider the definition of primitive gossip matrices.
Definition 2.1 (Primitive Matrix).
A matrix $W\in\mathbb{R}^{n\times n}$ is called primitive if there exists $k\in\mathbb{N}$ such that
$W^{k}>0\quad\text{(entrywise)}.$
Equivalently, a row-stochastic matrix is primitive if and only if the induced Markov chain is irreducible and aperiodic. In this case, $W$ has a unique stationary distribution $\pi>0$ with $\pi^{\top}W=\pi^{\top}$, and moreover
$W^{t}\to\mathbf{1}\pi^{\top}\qquad\text{as }t\to\infty.$
3 Privacy Analysis
3.1 Gossip Averaging as Discrete-Time Linear State-Space Dynamics
We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., antoulas2005approximation). In particular, we consider the systems without the feedthrough term in which case their dynamics is are described by the equations
$\theta_{t+1}$ $=A_{t}\theta_{t}+B_{t}u_{t},$
$y_{t}$ $=C_{t}\theta_{t},$
where at each time step $t$, $\theta_{t}$ represents the state vector, the matrix $A_{t}$ the state transition matrix, $B_{t}$ the input matrix which describes how the control input $u_{t}$ at time step $t$ affects the state, and the observation $y_{t}$ is related to the state vector $\theta_{t}$ by the observation matrix $C$.
We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{N}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[N]$ (see, e.g., Ch. 4 antoulas2005approximation). In our presentation, we focus on time-invariant graphs in which case $A_{t},B_{t}$ and $C_{t}$ are fixed.
3.2 Analysis of Non-Adaptive Compositions
We first show how to analyze DP gossip averaging with secure summation protocols for non-adaptive node-wise functions of data. To this end, without loss of generality, consider the task of globally averaging node-wise streams of data. I.e., each node $i$, $i\in[n]$, will have a stream $x_{0}^{i},x_{1}^{i},\dots$. Suppose each node adds normally distributed noise with variance $\sigma^{2}$ to its data point at each round. For $t\in[T]$, denote also $x_{t}=\begin{bmatrix}x_{t}^{1}&\ldots&x_{t}^{n}\end{bmatrix}^{\top}$. Denoting the global state variable at round $t$ with $\theta_{t}$, the gossip averaging with the gossip matrix $W$ can be written as
$\theta_{t+1}=W(\theta_{t}+x_{t}+u_{t}),$ (3.1)
where $u_{t}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. From the recursion (3.1) it follows that from the view determined by the selector $S$, the view is given by $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})=\begin{bmatrix}y_{1}&\ldots&y_{n}\end{bmatrix}^{\top}$, where
$y_{1}$ $=S(x_{0}+u_{0})$
$y_{2}$ $=S(Wx_{0}+Wu_{0}+x_{1}+u_{1})$
$y_{3}$ $=S(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1}+x_{2}+u_{2})$
$\vdots$
$y_{T-1}$ $=S(W^{T-1}x_{0}+\cdots+x_{T-1}+W^{T-1}u_{0}+\cdots+u_{T-1}),$
where $S$ is the selector matrix for $\mathcal{A}$
From this representation we get the following for a fixed pair of neighboring streams of data.
Lemma 7.
Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$. Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
$\mathcal{M}(D)=\widetilde{x}_{T}+H_{T}\widetilde{u}_{T}\quad\textrm{and}\quad\mathcal{M}(D^{\prime})=H_{T}\widetilde{u}_{T},$
where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T})$, and
$\widetilde{x}_{T}=H_{T}\left(\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}\otimes e_{j}\right)\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S&0&0&\dots&0\\ SW&S&0&\dots&0\\ SW^{2}&SW&S&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ SW^{T-1}&SW^{T-2}&\dots&SW&S\end{bmatrix}$ (3.2)
Applying the DP analysis of the projected Gaussian mechanism (Lemma 6) to the pair of distributions given by Lemma 7 directly gives the following result.
Theorem 8.
Consider the Gossip averaging algorithm (3.1) and suppose $D,D^{\prime}$ are neighboring datasets such that $D\simeq_{j}D^{\prime}$. Let $\widetilde{x}_{T}$ and $H_{T}$ be given as in Eq. (3.2). Let $H_{T}=U\Sigma V^{\top}$ be the compact SVD of $H_{T}$. Let $\widetilde{x}_{T}$ be of the form (3.2) and denote $c=\begin{bmatrix}c_{1}&c_{2}&\ldots&c_{T}\end{bmatrix}^{\top}\in\mathbb{R}^{T}$. Denote
$\Delta^{T}_{j\rightarrow i}=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2}=\max_{c\in\{-1,1\}^{T}}\left\lVert V^{\top}(c\otimes e_{j})\right\rVert_{2}\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|VV^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
where $\left|\cdot\right|$ denotes the element-wise absolute value of a matrix and $\mathbf{1}_{T}\in\mathbb{R}^{T}$ is the all-ones vector. Then, from the point-of-view of a node $i$, the node $j$ is tightly $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
3.3 Accounting for Nodes’ Knowledge of Injected Noise
We have so far assumed that the noise injected by the observing node (or subset of nodes) contributes to the DP guarantees of the other nodes. In practice, this likely has a small effect on the the DP guarantees. However, we can obtain rigorous guarantees by removing the noise terms corresponding to the observing nodes. This corresponds to removing suitable columns from the matrix $H_{T}$ appearing in the DP guarantees. Conditioning the privacy guarantees on the adversary’s knowledge about the randomness has been also considered recently by allouahprivacy is scheme called SecLDP.
For example, if the adversary is a single observing node and removes its own noise from the computations, instead of Thm. 8, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big(T\cdot(n-1)\big)$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq. (3.2) with the $i$th column vector of each $T\times n$-block column removed. Similar correction can be carried out also in all the other cases. We do this correction in all our experiments.
3.4 Adaptive Compositions and Connection to Matrix-Mechanism Accounting
The recent work by bellet2025unified shows that the linear systems perspective given in Theorem 8 can be equivalently analyzed via a so-called matrix mechanisms (kairouz2021practical; pillutla2025correlated). This analogy allows also extending the non-adaptive privacy analysis directly to adaptive privacy analysis. We next shortly discuss this connection. At a high level, both frameworks reduce an observer’s view to a linear image of a data-dependent signal plus correlated Gaussian noise.
Different Threat Scenarios.
The matrix mechanism formulation by bellet2025unified allows analyzing private decentralized optimization methods that have a linear structure. Similarly to our linear systems perspective, it also allows analyzing several existing threat models from the literature, namely:
  • The LDP approach where the adversary sees all messages (bellet2018personalized). This corresponds to the setting of Theorem 8 with $S=I_{n}$.
  • The pair-wise Nerwork Differential Privacy (PNDP) where the adversary is a node and see the messages it sends or receives (networkdp2). This corresponds to Theorem 8 with a selector matrix of a finite collection of nodes.
  • SecLDP (allouahprivacy), where the adversary can remove the randomization it contributes to the computations, thus weakening the privacy guarantees. This corresponds to deletion of certain columns in each block of the matrix $H_{T}$ given in Thmeorem 8, see Section 3.3.
Privacy Analysis of Matrix Mechanisms.
The matrix-mechanism analysis by bellet2025unified provides a unified viewpoint for characterizing privacy leakage in decentralized learning across different trust settings. The key observation is that the view of any adversarial coalition can be expressed as a linear measurement of a stacked vector of (possibly adaptive and time-varying) messages $G$, corrupted by a structured Gaussian noise term $Z$.
In the matrix mechanism formulation, the $d$-dimensional user-wise messages are stacked into a global matrix $G\in\mathbb{R}^{nT\times d}$ of the form
$G=\begin{bmatrix}G_{1}\\ \vdots\\ G_{T}\end{bmatrix}$,
where each block $G_{t}\in\mathbb{R}^{n\times d}$ corresponds to round $t$ and is given by
$G_{t}=\begin{bmatrix}g_{t}^{1}\\ \vdots\\ g_{t}^{n}\end{bmatrix}$,
with $g_{t}^{i}$ representing the message of node $i$ at iteration $t$. Thus, differences between neighboring datasets appear as differences along the rows $(g_{t}^{i})_{t=1}^{T}$ corresponding to a single user.
The privacy analyses of different private optimization methods under varying threat scenarios are then analyzed via the following results.
Theorem 9 (Bellet et al., 2025, Thm. 5).
For each of the three trust models, LDP, PNDP and SecLDP, and for every linear decentralized learning algorithm, there exist matrices $A$, $B$ and $C$ such that the adversarial view can be written as
$\mathcal{O}_{\mathcal{A}}=AG+BZ,$ (3.3)
with $A=BC$.
To quantify privacy, one needs to bound how much $G$ may vary under the adjacency relation while remaining observable through this linear mapping.
Definition 3.1 (Sensitivity under participation).
Let $B$ and $C$ be two matrices, and let $\Pi$ be a participation scheme for $G$. The sensitivity of $C$ with respect to $B$ under $\Pi$ is defined as
$\mathrm{sens}_{\Pi}(C;B):=\max_{G\simeq_{\Pi}G^{\prime}}\,\|\,C(G-G^{\prime})\,\|_{B^{+}B},$
where $\|v\|_{B^{+}B}:=\sqrt{v^{\top}(B^{+}B)v}$ denotes the seminorm induced by $B^{+}B$, and $G\simeq_{\Pi}G^{\prime}$ indicates that $G$ and $G^{\prime}$ differ only along entries that belong to the same individual according to $\Pi$.
A participation scheme $\Pi$ specifies which coordinates of the stacked update matrix $G$ may change under a neighboring-dataset relation. In other words, $\Pi$ encodes which entries of $G$ correspond to contributions of the same individual across time, ensuring that sensitivity is computed only over data differences permitted by the privacy definition.
Theorem 10 (Bellet et al., 2025, Thm. 7).
Let $\mathcal{O}_{\mathcal{A}}=AG+BZ$ be the attacker knowledge of a trust model on a linear decentralized learning algorithm, and denote by $\mathcal{M}(G)$ the corresponding mechanism. Let $\Pi$ be a participation scheme for $G$, and suppose $Z\sim\mathcal{N}(0,\nu^{2}I)$. Assume that $A$ is a column-echelon matrix and that $A=BC$ for some $C$. Then, if
$\nu=\sigma\,\mathrm{sens}_{\Pi}(C;B)\qquad\text{with}\qquad\mathrm{sens}_{\Pi}(C;B)\;\leq\;\max_{\pi\in\Pi}\;\sum_{s,t\in\pi}\bigl|\bigl(C^{\top}B^{+}BC\bigr)_{s,t}\bigr|,$
the mechanism $\mathcal{M}$ satisfies $\tfrac{1}{\sigma}$-Gaussian Differential Privacy, even when $G$ is chosen adaptively.
Adaptive Analysis.
Notably, bellet2025unified extend the existing adaptive analysis of matrix mechanisms (denisov2022improved) which apply to square coefficient matrices $A$ to the case where $A$ has less rows than columns. The crucial implication is that once the representation
$\mathcal{O}_{\mathcal{A}}=AG+BZ$
is identified, privacy guarantees for adaptive compositions are the same as in the case of non-adaptive compositions and directly follow from a single sensitivity bound on $(B,C)$ under the participation scheme $\Pi$.
Matrix Mechanism Formulation of Gossip Averaging.
We focus on the canonical case of gossip averaging of which DP analysis is equivalent to that of the Muffliato-SGD with a single Muffliato averaging step. As shown in (Appendix A.4 bellet2025unified), in this case the matrix mechanism representation is determined by the matrices $A=B=H_{T}$ and $C=I$, where $H_{T}$ is the matrix given in Thm. 8.
Equivalence of Sensitivity Bounds.
We also see from Thm. 10 that the sensitivity upper bound equals the upper bound given in our Thm. 8 in case the differing node participates in every round, since if $H_{T}=U\Sigma V^{T}$ is the compact SVD of $H_{T}$, $B=H_{T}$ and $C=I$, then
$\left|C^{\top}B^{+}BC\right|=\left|VV^{\top}\right|$
and the summation in the upper bound of Thm. 10 equals the upper bound given in our Thm. 8 in case the differing node participates in every round.
4 Sensitivity Growth and Asymptotic Rate for Secure Summation Protocol
We next state a result that shows that asymptotically, the privacy guarantees of gossip averaging with secure summation are similar as in case of central aggregation.
We first need the following auxiliary lemma.
Lemma 11.
Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic. Then there exist constants $M<\infty$ and $\rho\in(0,1)$ such that for all $t\geq 1$
$W^{t}\;=\;\Pi\;+\;R^{t},\qquad\Pi:=\mathbf{1}\,\pi^{\top},\qquad\|R^{t}\|\;\leq\;M\,\rho^{t},$
where $\pi^{\top}W=\pi^{\top}$, $\pi>0$, $\sum_{i}\pi_{i}=1$.
Remark 12.
A row-stochastic primitive gossip matrix $W$ has a unique eigenvalue $\lambda_{1}=1$ and the rest of the eigenvalues $\lambda_{2},\ldots,\lambda_{n}$ have absolute value < 1, and as the proof of Lemma 11 shows, $1-\rho$ equals the spectral gap $\lambda_{1}-\max_{i\geq 2}\left|\lambda_{i}\right|$.
Theorem 13.
Let $W\in\mathbb{R}^{n\times n}$ be primitive and row-stochastic with stationary distribution $\pi>0$, $\sum_{k=1}^{n}\pi_{k}=1$, $\pi^{\top}W=\pi^{\top}$. Let $H_{T}$ be the system matrix given in Lemma 7 with $S=e_{i}^{\top}$ (view of a single node, i.e., that of secure summation protocol) and define
$\Delta_{T}(j\to i):=\max_{c\in\{\pm 1\}^{T}}\left\|H_{T}^{+}H_{T}(c\otimes e_{j})\right\|_{2}$
as in Theorem 8. Then
$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
Remark 14.
Notice that if $W$ is doubly stochastic, $\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}=\frac{1}{n}$ and we asymptotically have the same sensitivity as in case of a central aggregator.
5 Experiments
Illustration of the DP Bounds for Secure Summation Protocols. We first consider numerical evaluation of the DP bounds we obtain from Thm. 8.
First, consider a synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, i.e., each of the $n$ users is connected to each other with probability $p$. We take the gossip matrix $W$ to be a row-stochastic matrix obtained from row-wise averaging. Second, we consider a network generated using a preferential-attachment model (barabasi1999emergence) with $n=200$ nodes, initialized from a fully connected core of 5 nodes. Each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees. The gossip matrix $W$ is obtained from row-wise averaging.
As Figure 1 illustrates, in case of the secure summation, the squared sensitivity $(\Delta^{T}(j\to i))^{2}$ for a randomly chosen $j$ grows asymptotically linearly w.r.t. $T$ as predicted by Thm. 13.
Figure 1: Left: Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.2$, Right: gossip matrix of a preferential-attachment model with $n=200$ nodes, and the scaled sensitivity $(\Delta^{T}(j\to i))^{2}/T$ as a function of $T$. We consider separately the cases when random nodes $i$ and $j$ are neighbors and not.
6 Conclusions
We presented a framework for analyzing privacy guarantees for DP gossip averaging that accounts for all injected noise, and showed that secure-summation-based averaging has asymptotic sensitivity comparable to central aggregation. Future work includes accelerating the linear-algebra routines needed to compute total sensitivities on large or sparse graphs and analyzing other decentralized methods with similar linear dynamics to assess their long-term privacy guarantees.
Appendix A Proof of Lemma 6
Proof.
Recall, the projected Gaussian mechanism is give as
$\mathcal{M}(D)=f(D)+AZ,$
where $f:\mathcal{D}\rightarrow\mathbb{R}^{m}$, $A\in\mathbb{R}^{m\times n}$ and $Z\sim\mathcal{N}(0,\sigma^{2}I_{n})$ for some $\sigma>0$.
Denote the compact SVD of $A$ as $A=U_{r}\Sigma_{r}V_{r}^{\top}$, where $r$ denotes the rank of $A$. Since the columns of $U_{r}$ give a basis for the subspace $\mathrm{Range(A)}$ and $U_{r}$ has orthonormal columns, $U_{r}U_{r}^{\top}$ gives a projector onto $\mathrm{Range(A)}$. Since $f(D)-f(D^{\prime})\in\mathrm{Range(A)}$, it holds that
$f(D)-f(D^{\prime})=U_{r}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big).$ (A.1)
By using the compact SVD of $A$, we have that for all $\alpha\geq 0$:
$H_{\alpha}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)$ $=H_{\alpha}\left(f(D)+AZ\,||\,f(D^{\prime})+AZ\right)$
$=H_{\alpha}\left(f(D)-f(D^{\prime})+AZ\,||\,AZ\right)$
$=H_{\alpha}\left(f(D)-f(D^{\prime})+U_{r}\Sigma_{r}V_{r}^{\top}Z\,||\,U_{r}\Sigma_{r}V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(U_{r}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)+U_{r}\Sigma_{r}V_{r}^{\top}Z\,||\,U_{r}\Sigma_{r}V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)+\Sigma_{r}V_{r}^{\top}Z\,||\,\Sigma_{r}V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)+V_{r}^{\top}Z\,||\,V_{r}^{\top}Z\right)$
[…]
where $\widetilde{Z}\sim\mathcal{N}(0,\sigma^{2}I_{r})$ and where before the third last equality we have carried out multiplication from the left by $U_{r}^{\top}$ and before the second last equality we have carried out multiplication from the left by $\Sigma_{r}^{-1}$. The last step follows from the fact that $V_{r}$ has orthonormal columns. We see that for all $\varepsilon\in\mathbb{R}$,
[…]
where the right-hand side gives the tight $\delta(\varepsilon)$ for the Gaussian mechanism with sensitivity $\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ and noise scale $\sigma$. Further, since $V_{r}$ has orthonormal columns,
$\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ $=\left\lVert V_{r}\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$
$=\left\lVert A^{+}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$
and the claim follows. ∎
Appendix B Proof of Lemma 7
Proof.
From the translational invariance of the hockey-stick divergence, we have
$H_{\alpha}\left(\mathrm{View}_{\mathcal{M}(D)}(\{i\})\,||\,\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\}))\right)$ $=H_{\alpha}\left(H_{T}\widehat{x}_{T}+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widehat{x}_{T}^{\prime}+H_{T}\widetilde{u}_{T}\right)$
$=H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right).$
By the compact SVD $H_{T}=U_{r}\Sigma_{r}V_{r}^{\top}$, where $r$ denotes the rank of $H_{T}$, we have that for all $\alpha\geq 0$,
$H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
$=$ $H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
$=$ $H_{\alpha}\left(\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
$=$ $H_{\alpha}\left(V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+V_{r}^{\top}\widetilde{u}_{T}\,||\,V_{r}^{\top}\widetilde{u}_{T}\right)$
[…]
where $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$. ∎
Appendix C Proof of Thm. 8
Proof.
For an upper bound of the sensitivity, we need to find the maximum of $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ when $\Delta x$ is in the convex and compact domain
$\mathcal{D}_{j}:=\{\alpha\otimes e_{j}:\alpha\in[-1,1]^{T}\}$
We see that maximizing $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is equivalent to finding the maximum of the quadratic form $(\Delta x)^{\top}P\Delta x$ in $\mathcal{D}_{j}$ where the projector matrix $P=V_{r}V_{r}^{\top}$ is positive semi-definite and a projection on to the row space of $H_{T}$. Thus, the optimum is found from somewhere at the boundary of $\mathcal{D}_{j}$, i.e., it holds that $\Delta x=\alpha\otimes e_{j}$, where $\alpha=\{-1,1\}^{T}$.
For the inequality, we use the trivial inequality
$\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}=(\Delta x)^{\top}V_{r}V_{r}^{\top}(\Delta x)\leq\mathbf{1}_{T}^{\top}\left|V_{r}V_{r}^{\top}\right|\mathbf{1}_{T}.$
Appendix D Proof of Lemma 11
Proof.
Since $W$ is primitive and row-stochastic, $1$ is a simple eigenvalue with right eigenvector $\mathbf{1}$ and left eigenvector $\pi^{\top}$, and all other eigenvalues satisfy $|\lambda|<1$ (Perron-Frobenius). Choose a basis matrix S=[ 1B]S=[\ $\mathbf${1}\ \ B\ ] for some $B\in\mathbb{R}^{n\times n-1}$ so that the first row of $S^{-1}$ is $\pi^{\top}$. Then
$S^{-1}WS\;=\;\begin{bmatrix}1&0\\ 0&T^{\prime}\end{bmatrix}$,
and taking powers gives
$W^{t}\;=\;S\begin{bmatrix}1&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\;=\;\underbrace{\mathbf{1}\,\pi^{\top}}_{=\ \Pi}\;+\;\underbrace{S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}}_{=:R^{t}}.$
To bound $\|R^{t}\|$, fix any matrix norm $\|\cdot\|$. Since $\rho(T^{\prime}):=\max\{|\lambda|:\lambda\in\mathrm{spec}(T^{\prime})\}<1$, the Gelfand formula implies that for any $\rho$ with $\rho(T^{\prime})<\rho<1$ there exists $C<\infty$ such that $\|(T^{\prime})^{t}\|\leq C\,\rho^{t}$ for all $t\geq 1$. Therefore
$\|R^{t}\|\;=\;\left\|S\begin{bmatrix}0&0\\ 0&(T^{\prime})^{t}\end{bmatrix}S^{-1}\right\|\;\leq\;\|S\|\,\|S^{-1}\|\,\|(T^{\prime})^{t}\|\;\leq\;\big(\|S\|\,\|S^{-1}\|\;C\big)\,\rho^{t}.$
Setting $M:=\|S\|\,\|S^{-1}\|\,C$ gives the claim. ∎
Appendix E Proof Theorem 13
Proof.
First, from the compact SVD we see that for any matrix $A$, the product $A^{+}A$ is the orthogonal projector onto $\mathrm{row}(A)$. Set $P_{T}:=H_{T}^{+}H_{T}$ to be the orthogonal projection onto row space of $H_{T}$. Let $c\in\{\pm 1\}^{T}$ and define $z:=c\otimes e_{j}\in\mathbb{R}^{nT}$. Our strategy is to study the sensitivity
$\Delta_{T}(c,j\to i)=\|P_{T}z\|_{2}$
and derive a bound for it that is independent of $c$.
Because $W$ is primitive and row-stochastic, by Lemma 11 there exist constants $M>0$ and $\rho\in(0,1)$ such that
$W^{t}=\Pi+R^{t},\qquad\Pi:=\mathbf{1}\pi^{\top},\qquad\|R^{t}\|_{2}\leq M\rho^{t}\quad\forall t\geq 0.$ (E.1)
The $(t,s)$ block of $H_{T}$ (with $0\leq s\leq t\leq T-1$) is given by
$H_{T}(t,s)=e_{i}^{\top}W^{t-s}=e_{i}^{\top}\Pi+e_{i}^{\top}R^{t-s}=\pi^{\top}+e_{i}^{\top}R^{t-s}.$
Define the stationary part of $H_{T}$ as $\bar{H}_{T}$ such that blocks of $\bar{H}_{T}$ are given as
$\bar{H}_{T}(t,s)=\pi^{\top}\quad\text{ for }s\leq t,\qquad\bar{H}_{T}(t,s)=0\quad\text{for }s>t.$
Furthermore, define the transient part
$K_{T}:=H_{T}-\bar{H}_{T}.$
Thus
$H_{T}=\bar{H}_{T}+K_{T}.$
We clearly have that
$\mathrm{row}(\bar{H}_{T})=\{(\alpha_{0}\pi,\alpha_{1}\pi,\dots,\alpha_{T-1}\pi)\in\mathbb{R}^{nT}:\alpha_{t}\in\mathbb{R}\}$
and we can take as an orthonormal basis for $\mathrm{row}(\bar{H}_{T})$ for example $\{e_{1}\otimes\pi/\|\pi\|_{2},\ldots,e_{T}\otimes\pi/\|\pi\|_{2}\}.$ Therefore the orthogonal projection onto $\mathrm{row}(\bar{H}_{T})$ is given by
$\bar{P}_{T}z=\sum_{\ell=1}^{T}\langle e_{\ell}\otimes\pi/\|\pi\|_{2},c\otimes e_{j}\rangle=c\;\otimes\;\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi.$
and we have that
$\|\bar{P}_{T}z\|_{2}^{2}=\sum_{t=0}^{T-1}\left\|c_{t}\,\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi\right\|_{2}^{2}=T\,\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
We next study the transient part and show using standard perturbation theory for projection matrices that the distance between the projectors $P_{T}$ and $\bar{P}_{T}$ satisfies
$\|P_{T}-\bar{P}_{T}\|_{2}=O(1/\sqrt{T}).$
This rate is derived using the Wedin theorem which states that the distance between the row spaces $\mathrm{row}(H_{T})$ and $\mathrm{row}(\bar{H}_{T})$ is a function of $\|K_{T}\|_{2}$ (since the perturbation $H_{T}-\bar{H}_{T}=K_{T}$) and the smallest non-zero singular value of $\bar{H}_{T}$, $\sigma_{\min}(\bar{H}_{T})$.
From the Toeplitz structure it follows that clearly $\|K_{T}\|_{2}$ equals the 2-norm of the first column block of $K_{T}$. Moreover, from the transient behavior stated in Eq. (E.1) and a geometric sum it follows that $\|K_{T}\|_{2}=O(1)$. We see that
$\bar{H}_{T}=U\otimes\pi^{\top}$
for
$U=\begin{pmatrix}1&0&\cdots&0\\ 1&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1\end{pmatrix}$.
The singular values of a Kronecker product $A\otimes B$ are the products of the singular values of $A$ and $B$. Thus, the non-zero singular values of $\bar{H}_{T}$ are:
$\sigma_{k}(\bar{H}_{T})=\sigma_{m}(U)\cdot\sigma_{l}(\pi^{\top}).$
Since $\pi^{\top}$ has rank $1$, $\sigma_{1}(\pi^{\top})=\|\pi^{\top}\|_{2}=\|\pi\|_{2}.$ We also know that the smallest singular value of $U$ is given by
$\sigma_{\min}(U)=\sqrt{T}.$
Thus, $\sigma_{\min}(\bar{H}_{T})=\sqrt{T}\|\pi\|_{2}$, and by the Wedin theorem we have that
$\|P_{T}-\bar{P}_{T}\|_{2}\leq\frac{2\|K_{T}\|_{2}}{\sigma_{\min}(\bar{H}_{T})}=O(1/\sqrt{T}).$
Let $\mathcal{E}_{T}=P_{T}-\bar{P}_{T}$, so $\|\mathcal{E}_{T}\|_{2}=O(1/\sqrt{T})$.
We expand $\|P_{T}z\|_{2}^{2}$ using $P_{T}=\bar{P}_{T}+\mathcal{E}_{T}$. Recall $z=c\otimes e_{j}$ and $\|z\|_{2}=\sqrt{T}$:
$\|P_{T}z\|_{2}^{2}=\|(\bar{P}_{T}+\mathcal{E}_{T})z\|_{2}^{2}=\|\bar{P}_{T}z\|_{2}^{2}+2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)+\|\mathcal{E}_{T}z\|_{2}^{2}.$ (E.2)
Using a norm bound on $\mathcal{E}_{T}$, we have
$|2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)|$ $\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}z\|_{2}$
$\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}\|_{2}\|z\|_{2}$
$\leq 2\left(\sqrt{T}\frac{\pi_{j}}{\|\pi\|_{2}}\right)\cdot O(1/\sqrt{T})\cdot\sqrt{T}=O(\sqrt{T}).$
Furthermore,
$\|\mathcal{E}_{T}z\|_{2}^{2}\leq\|\mathcal{E}_{T}\|_{2}^{2}\|z\|_{2}^{2}=O(1/T)\cdot T=O(1).$
Substituting these bounds back into Eq. (E.2):
$\|P_{T}z\|_{2}^{2}=T\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(\sqrt{T}).$
This holds for any choice of $c\in\{\pm 1\}^{T}$. The maximum is thus:
$(\Delta_{T}(j\to i))^{2}=\max_{c\in\{\pm 1\}^{T}}\|P_{T}z\|_{2}^{2}=T\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(\sqrt{T}).$
Dividing by $T$ and taking the limit:
$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\lim_{T\to\infty}\left(\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}+O(1/\sqrt{T})\right)=\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$