### v3

# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

Abstract
Achieving differential privacy (DP) guarantees in fully decentralized machine learning is challenging due to the absence of a central aggregator and varying trust assumptions among nodes. We present a framework for DP analysis of decentralized gossip-based averaging algorithms with additive node-level noise, from arbitrary views of nodes in a graph. We present an analytical framework based on a linear systems formulation that accurately characterizes privacy leakage between nodes. Our main contribution is showing that the DP guarantees are those of a Gaussian mechanism, where the growth of the squared sensitivity is asymptotically $O(T)$, where $T$ is the number of training rounds, similarly as in the case of central aggregation. As an application of the sensitivity analysis, we show that the excess risk of decentralized private learning for strongly convex losses is asymptotically similar as in centralized private learning.
1 Introduction
Common federated learning (FL) scenarios assume the presence of a central parameter server for coordinating model updates. In contrast, fully decentralized setups operate without a central orchestrator: compute nodes, each holding a private dataset, directly exchange model states or updates with a subset of peers. Such decentralized architectures offer advantages in scalability, fault tolerance, and robustness, but also introduce new algorithmic and privacy challenges.
Fully decentralized gradient-based optimization methods are typically distributed variants of gradient descent and can be broadly classified into two categories.
In random walk-based methods (Lopes and Sayed, 2007; Johansson et al., 2009; Mao et al., 2020), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix. The neighbor then updates its local parameters and passes the updated model to another randomly chosen neighbor. Over time, and under mild assumptions on the transition matrix and the network graph (e.g., connectedness, symmetry), the random walk ensures uniform coverage of all nodes. However, such sequential update dynamics inherently limits parallelism and scalability.
In gossip averaging-based methods (Boyd et al., 2006; Dimakis et al., 2010), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states. Each node updates its parameters by averaging over the received messages. These operations are repeated iteratively until the network reaches approximate consensus on the model parameters. Gossip protocols are naturally parallelizable and more scalable than random walk approaches, making them attractive for large-scale decentralized learning. This work focuses on such gossip-based protocols for DP decentralized learning.
DP in Centralized FL. In centralized FL (Kairouz et al., 2021c), distributed DP (Ullah et al., 2023; Kairouz et al., 2021a) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (Truex et al., 2019; Erlingsson et al., 2019; Bell et al., 2020). This not only eliminates the need to trust individual nodes but also enables lower per-user noise by aggregating over multiple contributions.
DP in Decentralized Optimization. The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (Dekker et al., 2025; Zhu et al., 2019; Mrini et al., 2024; Pasquini et al., 2023). Early decentralized DP methods achieve privacy by locally perturbing gradients or models (Huang et al., 2015; Bellet et al., 2018a; Xu et al., 2022), relying on local DP guarantees. This often imposes a poor utility-privacy tradeoff compared to centralized DP (Chan et al., 2012).
To address this, recent works propose pair-wise network DP (Cyffers et al., 2024, 2022; Cyffers and Bellet, 2022), where privacy loss is analyzed between pairs of nodes. These relaxations give better utility while accounting for the structure of decentralized communication. Cyffers et al. (2022) give Rényi differential privacy (RDP) bounds that take into account the limited view, however they exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of compositions (or training iterations in ML model training), making them unsuitable for practical ML model training with reasonable privacy guarantees. Recently, Bellet et al. (2025) considered the linear-systems viewpoint of decentralized optimization to cover adaptive compositions through the so called matrix mechanisms (Kairouz et al., 2021b; Pillutla et al., 2025) which leads to considerably more accurate privacy analysis than that of Cyffers et al. (2022). Nevertheless, as our asymptotic sensitivity result (our Theorem 13) shows, the full linear-systems framework is still a valuable tool for understanding the privacy-preserving properties of decentralized optimization.
Our Approach. In this work, we develop techniques to analyze network DP guarantees under various threat models, showing that the total sensitivity grows asymptotically as $O(\sqrt{T})$, resulting in asymptotic $O(T)$ growth of the RDP parameters, for example, representing a significant improvement over prior analyses. This network DP accounting is obtained by interpreting the dynamics of the gossip averaging as a linear state-space system and the view of individual nodes or subsets of nodes as projected Gaussian mechanisms (GMs). This also provides tools to analyze the network DP guarantees in different threat scenarios. In addition to including the threat model of (Cyffers et al., 2022), where the neighboring nodes see the plain messages sent by their neighbors, our linear systems view allows analyzing algorithms that incorporate a secure summation protocol between neighboring nodes.
Secure Aggregation in Decentralized Learning. While the privacy amplification effect coming from summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored. In current gossip-based systems (Cyffers et al., 2022), each node receives messages directly from its neighbors and can, in principle, view their full content. As we show, summation over a node’s neighborhood can give a meaningful privacy amplification. Several studies have proposed integrating MPC primitives into decentralized learning workflows (Jayaraman et al., 2018; Lian et al., 2018; Jeon et al., 2021). Concrete protocols for decentralized secure aggregation have been developed in recent works (Sabater et al., 2022; Pereira et al., 2024; Biswas et al., 2024). The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (Nicolas et al., 2024).
Our Contributions:
  • We provide an analytical framework for evaluating differential privacy guarantees in decentralized gossip-based averaging algorithms. Our framework leverages a linear dynamical systems perspective to track sensitivity propagation through iterative updates.
  • We demonstrate that the DP guarantees can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed via the gossip matrix and the linear systems perspective we consider.
  • We show that in case of a single node view (e.g., that of secure summation based averaging protocols), the squared sensitivity scales asymptotically as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on RDP (Cyffers et al., 2022).
  • As an application of the sensitivity analysis, we combine the privacy bounds with the utility analysis of noisy decentralized learning (Koloskova et al., 2020) and show that the excess risk of strongly convex losses is asymptotically similar to the centralized private learning.
Relation to Recent Work by (Bellet et al., 2025). Recent work by (Bellet et al., 2025) builds on the linear systems formulation of private gossip algorithms and further represents it as a matrix mechanism (Kairouz et al., 2021b; Pillutla et al., 2025). Using this formulation, they show that the privacy guarantees extend to adaptive compositions, which we also rely on in this work.
Our contributions are complementary. While (Bellet et al., 2025) focus on adaptive privacy and general sensitivity bounds via the matrix mechanism view, we work directly in the linear systems formulation to derive analytic characterizations of both privacy and convergence for decentralized learning, which require this representation and are not obtained in their analysis. Our results show that private decentralized learning can match centralized methods in both privacy and utility.
2 Background
We first shortly review the required technicalities on differential privacy. We then discuss the projected Gaussian mechanism central to our analysis and define the model for decentralized learning and the network DP guarantees.
2.1 Differential Privacy
An input dataset containing $n$ data points is denoted as $D=(x_{1},\ldots,x_{n})\in\mathcal{D}$, where $\mathcal{D}$ denotes the set of datasets of all sizes. We say that two datasets $D$ and $D^{\prime}$ are neighbors if we get one by adding or removing one element to/from the other (denoted $D\sim D^{\prime}$). We say that a mechanism $\mathcal{M}\,:\,\mathcal{D}\rightarrow\mathcal{O}$ is $(\varepsilon,\delta)$-DP if the output distributions for neighboring datasets are always $(\varepsilon,\delta)$-indistinguishable.
Definition 1 (Dwork et al. 2006).
Let $\varepsilon\geq 0$ and $\delta\in[0,1]$. A randomized mechanism $\mathcal{M}\,:\,\mathcal{D}\rightarrow\mathcal{P}(\mathcal{O})$, where $\mathcal{P}(\mathcal{O})$ denotes the set of probability distributions over the output space $\mathcal{O}$, is $(\varepsilon,\delta)$-DP if for every pair of neighboring datasets $D,D^{\prime}$, every measurable set $O\subset\mathcal{O}$,
$\mathbb{P}(\mathcal{M}(D)\in O)\leq{\rm e}\hskip 1.0pt^{\epsilon}\mathbb{P}(\mathcal{M}(D^{\prime})\in O)+\delta.$
Given two distributions $(P,Q)$, if the conditions $\mathbb{P}(P\in O)\leq{\rm e}\hskip 1.0pt^{\epsilon}\mathbb{P}(Q\in O)+\delta$ and $\mathbb{P}(Q\in O)\leq{\rm e}\hskip 1.0pt^{\epsilon}\mathbb{P}(P\in O)+\delta$ hold for every measurable set $O\subset\mathcal{O}$, we also denote $P\simeq_{(\varepsilon,\delta)}Q$.
With the hockey-stick divergence we can equivalently measure the $(\varepsilon,\delta)$-distance of distributions. For probability distributions $P$ and $Q$, and for $\varepsilon\in\mathbb{R}$, it is defined as $H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(P||Q)=\int[P(t)-{\rm e}\hskip 1.0pt^{\varepsilon}Q(t)]_{+}$, where $z]_{+}=\max\{0,z\}$. The $(\varepsilon,\delta)$-DP guarantees can be then given as follows.
Lemma 2 (Balle et al. 2018).
A mechanism $\mathcal{M}$ satisfies $(\epsilon,\delta)$-DP if and only if, $\max_{D\sim D^{\prime}}H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(\mathcal{M}(X)||\mathcal{M}(X^{\prime}))\leq\delta$.
The Gaussian Mechanism is a common way to achieve $(\epsilon,\delta)$-differential privacy by adding Gaussian noise to a functions output.
Definition 3 (Gaussian Mechanism).
Let $f:\mathcal{D}\to\mathbb{R}^{d}$ be a function with $\ell_{2}$-sensitivity defined as $\Delta_{2}(f)=\max_{D,D^{\prime}}\|f(D)-f(D^{\prime})\|_{2}$, where the maximum is over all adjacent datasets $D,D^{\prime}$ (i.e., datasets differing in at most one individual’s data). The Gaussian mechanism outputs
$\mathcal{M}(D)=f(D)+Z,$
where $Z\sim\mathcal{N}(0,\sigma^{2}I_{d})$.
Running a DP training algorithm for $T$ iterations is commonly modeled as an adaptive composition of $T$ mechanisms such that the adversary has a view on the output of all intermediate outputs. This means that we then analyze mechanisms of the form
$\mathcal{M}^{(T)}(D)=\big(\mathcal{M}_{1}(D),\mathcal{M}_{2}(\mathcal{M}_{1}(D),D),\ldots,\mathcal{M}_{T}(\mathcal{M}_{1}(D),\ldots,\mathcal{M}_{T-1}(D),D)\big).$
The results of (Balle and Wang, 2018) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism. Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by Dong et al. (2022).
Lemma 4 (Dong et al. 2022).
Consider an adaptive composition of $T$ Gaussian mechanisms, each with $L_{2}$-sensitivity $\Delta$ and noise scale parameter $\sigma$. The adaptive composition is $(\varepsilon,\delta)$-DP for
$\delta(\varepsilon)=\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}\cdot\Delta}+\frac{\sqrt{T}\cdot\Delta}{2\sigma}\right)-e^{\varepsilon}\Phi\left(-\frac{\varepsilon\sigma}{\sqrt{T}\cdot\Delta}-\frac{\sqrt{T}\cdot\Delta}{2\sigma}\right).$
An alternative and often more convenient way to analyze privacy guarantees is via Rényi Differential Privacy (RDP) (Mironov, 2017), which is based on Rényi divergence. For $\alpha>1$, the Rényi divergence of order $\alpha$ between two distributions $P$ and $Q$ is defined as
$D_{\alpha}(P\|Q)=\frac{1}{\alpha-1}\log\mathbb{E}_{Q}\!\left[\left(\frac{P}{Q}\right)^{\alpha}\right].$
A mechanism $\mathcal{M}$ is said to satisfy $(\alpha,\varepsilon)$-RDP if for all neighboring datasets $D\sim D^{\prime}$, $D_{\alpha}(\mathcal{M}(D)\|\mathcal{M}(D^{\prime}))\leq\varepsilon.$ The Gaussian mechanism with $\ell_{2}$-sensitivity $\Delta$ and noise variance $\sigma^{2}$ satisfies $(\alpha,\varepsilon)$-RDP with $\varepsilon=\frac{\alpha\,\Delta^{2}}{2\sigma^{2}}.$
We next give a result for a certain projected Gaussian mechanism, where the noise vector is projected to a smaller-dimensional subspace.
2.2 Projected Gaussian Mechanism and Moore–Penrose Pseudoinverse
In our results, the network DP guarantees become those of a Gaussian mechanism of the form
$\mathcal{M}(D)=f(D)+AZ,$ (2.1)
where $f:\mathcal{D}\rightarrow\mathbb{R}^{m}$, $A\in\mathbb{R}^{m\times n}$ and $Z\sim\mathcal{N}(0,\sigma^{2}I_{n})$ for some $\sigma>0$. To analyze the DP guarantees of these mechanism, we will use the Moore–Penrose pseudoinverse of $A$. We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (Golub and Van Loan, 2013).
Definition 5 (Compact SVD and Moore–Penrose Pseudoinverse).
Let $A\in\mathbb{R}^{m\times n}$ be a matrix of rank $r$. The compact SVD (also known as the economy SVD) of $A$ is given by:
$A=U_{r}\Sigma_{r}V_{r}^{\top}$
where $U_{r}\in\mathbb{R}^{m\times r}$ and $V_{r}\in\mathbb{R}^{n\times r}$ contain the left and right singular vectors corresponding to the non-zero singular values, respectively, and $\Sigma_{r}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the non-zero singular values $\sigma_{1}\geq\dots\geq\sigma_{r}>0$. The Moore–Penrose pseudoinverse of $A$, denoted $A^{+}$, is defined via compact SVD as
$A^{+}=V_{r}\Sigma_{r}^{+}U_{r}^{\top},$
where $\Sigma_{r}^{+}\in\mathbb{R}^{r\times r}$ is a diagonal matrix with entries $\frac{1}{\sigma_{1}},\dots,\frac{1}{\sigma_{r}}$, i.e., $\Sigma_{r}^{+}=\mathrm{diag}\left(\frac{1}{\sigma_{1}},\frac{1}{\sigma_{2}},\dots,\frac{1}{\sigma_{r}}\right).$
To analyze the gossip averaging algorithms, the following technical result will play a central role.
Lemma 6.
Let $\sigma>0$ and $A\in\mathbb{R}^{m\times n}$. Suppose $\mathcal{M}(D)$ is a projected Gaussian mechanism of the form given in Eq. (2.1). Let $D$ and $D^{\prime}$ be two datasets such that $f(D)-f(D^{\prime})\in\mathrm{Range(A)}$, where $\mathrm{Range(A)}$ denotes the subspace spanned by the columns of $A$. Then, for all $\alpha\geq 0$,
$H_{\alpha}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)=H_{\alpha}\left(\mathcal{N}\big(\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2},\sigma^{2}\big)\,||\,\mathcal{N}(0,\sigma^{2})\right).$
i.e., the $(\varepsilon,\delta)$-distance between $\mathcal{M}(D)$ and $\mathcal{M}(D^{\prime})$ is upper bounded by the $(\varepsilon,\delta)$-DP guarantee of a gaussian mechanism with $L_{2}$-sensitivity $\left\lVert A^{+}(f(D)-f(D^{\prime}))\right\rVert_{2}$ and noise parameter $\sigma$.
2.3 Decentralized Learning and Network Differential Privacy
We consider a connected undirected graph $G=(V,E)$ with $|V|=n$ nodes. The graph structure is encoded by an $n\times n$ boolean adjacency matrix and we denote the nodes by indices $n]=\{1,\dots,n\}$. For each node $i\in[n]$, the neighborhood and closed neighborhood are defined as
$N_{i}=\{j\in[n]:(i,j)\in E\},\qquad\bar{N}_{i}=N_{i}\cup\{i\}.$
Local Data and Learning Objective. The global dataset is partitioned across nodes:
$D=\bigcup_{i=1}^{n}D_{i},$
where node $i$ holds the shard $D_{i}$. The goal of decentralized learning is to minimize the global loss
$f(\theta)=\sum_{i=1}^{n}f_{i}(\theta),\qquad\theta\in\mathbb{R}^{d},$
where $f_{i}(\theta)$ depends only on $D_{i}$. At each round $t\in[T]$, node $i$ maintains a state variable $\theta_{t}]_{i}$. For the analysis, we consider the scalar case ($d=1$); the multidimensional case follows from the matrix mechanisms analogy given by Bellet et al. (2025) as described in Section A.
Dataset Adjacency. We adopt a record-level adjacency relation. Two datasets $D=\cup_{i}D_{i}$ and $D^{\prime}=\cup_{i}D_{i}^{\prime}$ are adjacent, written $D\sim D^{\prime}$, if they differ in exactly one node’s data:
$D\sim_{i}D^{\prime}\quad\iff\quad D_{j}=D^{\prime}_{j}\ \text{for all }j\neq i.$
Mechanism and View. We define a decentralized mechanism as a randomized mapping
$\mathcal{M}:D\mapsto\text{(all variables exchanged during the protocol)}.$
For a set of nodes $\mathcal{A}\subseteq[n]$, the view of $\mathcal{M}$ from the point-of-view of $\mathcal{A}$, i.e., the part of $\mathcal{M}(D)$ seen by nodes in $\mathcal{A}$, is denoted by
$\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A}).$
To extract node-wise components, we define the selector matrix: for $\mathcal{A}=\{i_{1},\dots,i_{|\mathcal{A}|}\}$,
$S(\mathcal{A})=\sum_{k=1}^{|\mathcal{A}|}e_{k}e_{i_{k}}^{\top}\in\mathbb{R}^{|\mathcal{A}|\times n},$
where $e_{k}$ is the $k$th standard basis vector.
Network Differential Privacy. Node $j$ is $(\varepsilon,\delta)$-DP from the point of view of $\mathcal{A}\subseteq[n]$ if for all adjacent datasets $D\sim_{j}D^{\prime}$,
$\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})\;\simeq_{(\varepsilon,\delta)}\;\mathrm{View}_{\mathcal{M}(D^{\prime})}(\mathcal{A}).$
We assume honest-but-curious nodes: they follow the protocol but analyze observed messages. As in (Cyffers et al., 2022), we first allow noise added by observers to contribute to privacy; this assumption is removed in Section 3.3.
Gossip-Based Averaging. Let $x_{t}\in\mathbb{R}^{n}$ denote the vector of local values exchanged at time $t$. A gossip update is given by
$\theta_{t+1}=Wx_{t},$
where $W\in\mathbb{R}^{n\times n}$ is the gossip matrix. We call $W$:
$\text{row-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1},$
and
$\text{doubly-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1}\ \text{and}\ W^{\top}\mathbf{1}=\mathbf{1}.$
Primitive gossip matrices have vanishing transient components, and we focus on this class in our asymptotic analysis.
Definition 7 (Primitive Matrix).
A matrix $W\in\mathbb{R}^{n\times n}$ is called primitive if there exists $k\in\mathbb{N}$ such that
$W^{k}>0\quad\text{(entry-wise)}.$
Equivalently, a row-stochastic matrix is primitive if and only if the induced Markov chain is irreducible and aperiodic. In this case, $W$ has a unique stationary distribution $\pi>0$ with $\pi^{\top}W=\pi^{\top}$, and moreover
$W^{t}\to\mathbf{1}\pi^{\top}\qquad\text{as }t\to\infty.$
Remark 8.
A primitive gossip matrix $W$ has a unique eigenvalue $\lambda_{1}=1$ and the rest of the eigenvalues $\lambda_{2},\ldots,\lambda_{n}$ have absolute value strictly less than 1.
We repeatedly use the following auxiliary lemma to bound the transient part of $W$.
Lemma 9.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic, and primitive. Then there exists $\rho\in(0,1)$ and a symmetric matrix $R$ such that
$W=\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}+R,$
$\mathbf{1}_{n}^{\top}R=0$, and for all integers $t\geq 0$, $\|R^{t}\|_{2}\leq\rho^{t}.$
3 Privacy Analysis
3.1 Gossip Averaging as Discrete-Time Linear State-Space Dynamics
We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., Antoulas, 2005). In particular, we consider the systems without the feedthrough term in which case their dynamics are described by the equations
$\theta_{t+1}$ $=A_{t}\theta_{t}+B_{t}u_{t},$
$y_{t}$ $=C_{t}\theta_{t},$
where at each time step $t$, $\theta_{t}$ represents the state vector, the matrix $A_{t}$ the state transition matrix, $B_{t}$ the input matrix which describes how the control input $u_{t}$ at time step $t$ affects the state, and the observation $y_{t}$ is related to the state vector $\theta_{t}$ by the observation matrix $C$.
We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{T}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[T]$ (see, e.g., Ch. 4 Antoulas, 2005). In our presentation, we focus on time-invariant graphs in which case $A_{t},B_{t}$ and $C_{t}$ are fixed.
3.2 DP Analysis of Compositions
We show how to analyze DP gossip averaging algorithms using the linear systems view. We first focus on non-adaptive compositions, where the node-wise contributions do not depend on the state variables. The connection between the linear systems view and matrix mechanisms as established by Bellet et al. (2025) then allows extending all the results to the adaptive case. To this end, without loss of generality, consider the task of globally averaging node-wise streams of data. I.e., each node $i$, $i\in[n]$, will have a stream $x_{0}^{i},x_{1}^{i},\dots$. Suppose each node adds normally distributed noise with variance $\sigma^{2}$ to its data point at each round. For $t\in[T]$, denote also $x_{t}=\begin{bmatrix}x_{t}^{1}&\ldots&x_{t}^{n}\end{bmatrix}^{\top}$. Denoting the global state variable at round $t$ with $\theta_{t}$, the gossip averaging with the gossip matrix $W$ can be written as
$\theta_{t+1}=W(\theta_{t}+x_{t}+u_{t}),$ (3.1)
where $u_{t}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. From the recursion (3.1) it follows that from the view determined by the selector $S$, the view is given by $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})=\begin{bmatrix}y_{1}&\ldots&y_{n}\end{bmatrix}^{\top}$, where
$y_{1}$ $=S(x_{0}+u_{0})$
$y_{2}$ $=S(Wx_{0}+Wu_{0}+x_{1}+u_{1})$
$y_{3}$ $=S(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1}+x_{2}+u_{2})$
$\vdots$
$y_{T-1}$ $=S(W^{T-1}x_{0}+\cdots+x_{T-1}+W^{T-1}u_{0}+\cdots+u_{T-1}),$
where $S$ is the selector matrix for $\mathcal{A}$
From this representation we get the following for a fixed pair of neighboring streams of data.
Lemma 10.
Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $c_{t}\in\mathbb{R}$. Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of the node set $\mathcal{A}\subset[n]$, where $(\varepsilon,\delta)$ is given by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
$\mathcal{M}(D)=\widetilde{x}_{T}+H_{T}\widetilde{u}_{T}\quad\textrm{and}\quad\mathcal{M}(D^{\prime})=H_{T}\widetilde{u}_{T},$
where $\widetilde{u}_{T}\sim\mathcal{N}(0,\sigma^{2}I_{T})$, and
$\widetilde{x}_{T}=H_{T}\left(\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}\otimes e_{j}\right)$
and
$H_{T}:=\begin{bmatrix}S&0&0&\dots&0\\ SW&S&0&\dots&0\\ SW^{2}&SW&S&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ SW^{T-1}&SW^{T-2}&\dots&SW&S\end{bmatrix}$, (3.2)
and $S$ is the selector matrix for $\mathcal{A}$.
Remark 11.
For example, in case we consider the threat model where all the messages sent by the neighbors are visible to a given node $i$, the selector matrix $S$ can be taken as $S=S(\bar{N}_{i})$ corresponding to the view of a closed neighborhood of node $i$. In case of secure summation over the nearest neighbors, we may take $S=e_{i}^{\top}$, i.e., the view of a single node.
Applying the DP analysis of the projected Gaussian mechanism (Lemma 6) to the pair of distributions given by Lemma 10 gives the following general $(\varepsilon,\delta)$-DP bound the noisy gossip averaging.
Theorem 12.
Consider the Gossip averaging algorithm (3.1) and suppose $D,D^{\prime}$ are neighboring datasets such that $D\simeq_{j}D^{\prime}$. Suppose also that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$. Let $\widetilde{x}_{T}$ and $H_{T}$ be given as in Eq. (3.2) for some $\mathcal{A}\subset[n]$. Let $H_{T}=U\Sigma V^{\top}$ be the compact SVD of $H_{T}$. Let $\widetilde{x}_{T}$ be of the form given in Eq. (3.2). Denote
$\Delta^{T}_{j\rightarrow\mathcal{A}}$ $=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2}$
$=\max_{c\in\{-1,1\}^{T}}\left\lVert V^{\top}(c\otimes e_{j})\right\rVert_{2}$
$\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|VV^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
where $\left|\cdot\right|$ denotes the element-wise absolute value of a matrix and $\mathbf{1}_{T}\in\mathbb{R}^{T}$ is the all-ones vector. Then, from the point-of-view of the node set $\mathcal{A}$, the node $j$ is $(\varepsilon,\delta)$-DP, where $(\varepsilon,\delta)$ is the privacy guarantee of the Gaussian mechanism with sensitivity $\Delta^{T}_{j\rightarrow i}$ and noise scale $\sigma$.
As we outline in Appendix A, due to the connection between the linear systems view and the so called matrix mechanisms, as shown by Bellet et al. (2025), Theorem 12 holds also in case adaptive cade, i.e., when the node-wise contributions $x_{t}$ depend on the state variable $\theta_{t}$. Moreover, from the spherical symmetry of the Gaussian noise it follows that the results generalize to arbitrary dimensions, when the condition $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$, is replaced by 2-norm bound for the sensitivity of node $j$’s contribution. We also remark that the sensitivity $\Delta^{T}_{j\rightarrow\mathcal{A}}$ leads to a tight $(\varepsilon,\delta)$-DP bound.
3.3 Accounting for Nodes’ Knowledge of Injected Noise
We have so far assumed that the noise injected by the nodes belonging to the observing set $\mathcal{A}\in[n]$ contributes to the DP guarantees of the other nodes. We can obtain rigorous guarantees also in case they do not contribute, by removing the noise terms corresponding to the observing nodes. This corresponds to removing suitable columns from the matrix $H_{T}$ appearing in the DP guarantees. Conditioning the privacy guarantees on the adversary’s knowledge about the randomness has been also considered recently by Allouah et al. (2024) is scheme called SecLDP.
For example, if the adversary is a single observing node and removes its own noise from the computations, instead of Thm. 12, we get the DP guarantees for node $j$ from the point of view node $i$ using the sensitivity $\Delta^{T}_{j\rightarrow i}=\left\lVert\widehat{H}_{T}^{+}\widetilde{x}_{T}\right\rVert_{2}$ where $\widehat{H}_{T}$ is the $T\times\big(T\cdot(n-1)\big)$ matrix corresponding to the $T\times(T\cdot n)$ matrix $H_{T}$ of Eq. (3.2) with the $i$th column vector of each $T\times n$-block column removed. For a general adversarial set $\mathcal{A}$, we obtain a complete analogy of Lemma 10 and Thm. 12 with the matrix blocks $SW^{k}$, $k=0,1,\ldots,T-1$, replaced by $SW^{k}S_{2}$, where $S_{2}$ selects the columns corresponding to the indices in the complement of $\mathcal{A}$.
3.4 Compute Efficient Upper Bound
We obtain a compute efficient upper bound for the sensitivity from a quadratic-form representation. Since $H_{T}$ has full row rank, $K_{T}:=H_{T}H_{T}^{\top}$ is invertible and for all $v\in\mathbb{R}^{m\cdot T\times 1}$
$\left\|H_{T}^{+}H_{T}v\right\|_{2}^{2}=(H_{T}v)^{\top}K_{T}^{-1}(H_{T}v).$
In particular, for $c\in\{-1,1\}^{T}$ and $v(c)=c\otimes e_{j}$, letting $b(c):=H_{T}v(c)$ we have
$\Delta_{T}(j\to\mathcal{A})^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}K_{T}^{-1}b(c).$
Moreover, we can construct a matrix $G_{T}\in\mathbb{R}^{mT\times T}$ such that $b(c)=G_{T}c.$ Hence, we can construct an $Tm\times Tm$ matrix $M_{T}:=G_{T}^{\top}K_{T}^{-1}G_{T}\succeq 0$, with which we get the efficiently computable upper bound
$\Delta_{T}(j\to\mathcal{A})^{2}$ $=\max_{c\in\{\pm 1\}^{T}}c^{\top}M_{T}c$
$\leq\max_{\|x\|_{2}^{2}=T}x^{\top}M_{T}x$
$=T\,\lambda_{\max}(M_{T}).$
We illustrate this bound numerically in Appendix F.
4 Asymptotic Sensitivity Growth Rate
We next state a result that shows that asymptotically, the privacy guarantees of gossip averaging with secure summation are similar as in case of central aggregation.
Theorem 13.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic and primitive. Let $H_{T}$ be the system matrix given in Lemma 10 with $\mathcal{A}=\{i\}$, i.e., $S=e_{i}^{\top}$ for some $i\in[T]$ (single-node view) and define
$\Delta_{T}(j\to i):=\max_{c\in\{\pm 1\}^{T}}\left\|H_{T}^{+}H_{T}(c\otimes e_{j})\right\|_{2}$
as in Theorem 12. Then
$\lim_{T\to\infty}\frac{(\Delta_{T}(j\to i))^{2}}{T}=\frac{1}{n}.$
Remark 14.
Notice that $(\Delta_{T}(j\to i))^{2}\sim T/n$ matches the sensitivity of the centralized case when each of the $n$ devices adds independent Gaussian noise with variance $\sigma^{2}$: their average has the output noise variance $n\sigma^{2}$ and as a result the DP analysis is equivalent to that of a Gaussian mechanism with noise scale $\sigma$ and squared sensitivity $T/n$ over a $T$-wise composition.
Using the proof technique of Thm. 13, we can also obtain the following sensitivity bound for the case the observing set $\mathcal{A}$ is of size $m$, $m>1$. Notice that Theorem 15 does not give a lower bound: as the proof indicates, the tight sensitivity becomes dependent on the gossip matrix $W$ and the observing set $\mathcal{A}$. Although the proof of Theorems 15 and 16 are analogous to the proof of Theorem 13 we state all these results separately for readability.
Theorem 15.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly-stochastic, and primitive. Fix a set of observing nodes $\mathcal{A}\subseteq\{1,\dots,n\}$ with $m:=|\mathcal{A}|$, and let $S_{\mathcal{A}}\in\mathbb{R}^{m\times n}$ be the row-selector matrix that extracts the coordinates indexed by $\mathcal{A}$ (i.e., $S_{\mathcal{A}}x=x_{\mathcal{A}}$). Let $H_{T}$ be the system matrix of Lemma 10 (Eq. (3.2)) constructed with observation matrix $S=S_{\mathcal{A}}$.
For any $j\in\{1,\dots,n\}$ define the sensitivity
$\Delta_{T}(j\to\mathcal{A}):=\max_{c\in\{\pm 1\}^{T}}\Bigl\|H_{T}^{+}H_{T}\,(c\otimes e_{j})\Bigr\|_{2}.$
Then, denoting $\rho:=\max\{\left|\lambda_{2}(W)\right|,\left|\lambda_{n}(W)\right|\}$, we have that for all $T\geq 1$,
$\frac{\Delta_{T}(j\to\mathcal{A})^{2}}{T}\leq\frac{m}{n}\,T+\frac{2}{(1-\rho)^{2}}\sqrt{\frac{m}{n}\,T}+\frac{2}{(1-\rho)^{4}}$
which further implies $\lim_{T\to\infty}\frac{\Delta_{T}(j\to\mathcal{A})^{2}}{T}\leq\frac{m}{n}\,T$.
Using the proof technique of Thm. 13, we also get an asymptotic upper bound for the case there are $m$ observing nodes and the noise injected by them does not contribute to the DP guarantees. We remark that the proof of Thm. 16 gives a similar explicit upper bound as in Thm. 15.
Theorem 16.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic, and primitive. Fix an observing set $\mathcal{A}\subseteq\{1,\dots,n\}$ with $m:=|\mathcal{A}|$. Suppose $j\notin\mathcal{A}$ and suppose the noise terms injected by the nodes in $\mathcal{A}$ do not contribute to the DP guarantees. Denote by $\bar{\Delta}_{T}(j\to\mathcal{A})$ the sensitivity of Theorem 12 with the blocks $SW^{k}$, $k=0,1,\ldots,T-1$, replaced by $SW^{k}S_{2}$, where $S_{2}$ selects the columns corresponding to the indices in the complement of $\mathcal{A}$. Then,
$\limsup_{T\to\infty}\frac{\bar{\Delta}_{T}(j\to\mathcal{A})^{2}}{T}\leq\frac{m}{n-m}.$
4.1 Experimental Illustration of the DP Bounds for Single Node View
We next consider a numerical illustration of the bound given by Theorem 13 in case a single node is observed. First, consider a randomly drawn synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$, i.e., each of the $n$ users is connected to each other with probability $p$. Second, consider a randomly drawn graph generated using a preferential-attachment model (Barabási and Albert, 1999) with $n=100$ nodes, initialized from a fully connected core of 5 nodes. Each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees. This model produces a structure characterized by heavy-tailed degree distributions. In both cases, we take the gossip matrix $W$ to be a doubly stochastic matrix obtained from so called Metropolis–Hastings averaging of a given symmetric adjacency matrix.
As Figure 1 illustrates, in case only a single node $i$ is viewed, the squared sensitivity $(\Delta^{T}(j\to i))^{2}$ grows asymptotically linearly w.r.t. $T$ with the rate $1/n$ as predicted by Thm. 13. Here both $i$ and $j$ are chosen randomly.
Figure 1: The scaled sensitivity $(\Delta_{T}(j\to i))^{2}/T$ as a function of $T$ for a) randomly drawn Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$ and b) a randomly drawn doubly-stochastic gossip matrix of a preferential-attachment model with $n=100$ nodes. We consider separately the cases when random nodes $i$ and $j$ are neighbors and not.
5 Utility Bound for Strongly Convex Losses
We next show how our sensitivity analysis can be plugged into standard convergence analysis of decentralized optimization (Koloskova et al., 2020) to obtain explicit privacy-utility bounds for private decentralized learning based on gossip averaging. For simplicity, we focus on the case of single-node view, however the results can be directly generalized to multi-node view using Theorems 15 and 16.
Setup. We consider decentralized minimization of the loss over $n$ nodes,
$f(\theta)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$
over a connected undirected graph with a fixed symmetric, doubly-stochastic gossip matrix $W$. Assume each $f_{i}$ is $L$-smooth and $\mu$-strongly convex. Each node performs full-batch DP gradient descent, followed by gossip averaging:
$\theta_{i}^{t+\frac{1}{2}}$ $=\theta_{i}^{t}-\eta\,(\nabla f_{i}(\theta_{i}^{t})+\xi_{i}^{t}),$ (5.1)
$\theta_{i}^{t+1}$ $=\sum_{j=1}^{n}w_{ij}\,\theta_{j}^{t+\frac{1}{2}},$
where $\xi_{i}^{t}\sim\mathcal{N}(0,\sigma_{\mathrm{DP}}^{2}I_{d})$ are independent across $i,t$.
Following (Koloskova et al., 2020), we measure utility in terms of the average iterate $\bar{\theta}^{t}=\frac{1}{n}\sum_{i=1}^{n}\theta_{i}^{t}.$
Topology parameter (absolute spectral gap). Let $1=\lambda_{1}(W)\geq\lambda_{2}(W)\geq\cdots\geq\lambda_{n}(W)\geq-1$ be eigenvalues of $W$ and define the absolute spectral gap
$\gamma:=1-\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}\in(0,1].$ (5.2)
In the decentralized SGD analysis by Koloskova et al. (2020), the effect of the network enters through a consensus-rate parameter $p$; for fixed symmetric $W$ one may take (see Lemma J.1 in Appendix)
$p=1-\gamma^{2}\;\;\;\text{and hence}\;\;\;\frac{1}{p}=O\!\left(\frac{1}{\gamma}\right).$ (5.3)
Using Sensitivity Bound to Set DP Noise. Under the threat model of a single-node view $S=e_{i}^{\top}$, our Theorem 13 shows that the total sensitivity from node $j$ to observer $i$, $\Delta_{T}(j\to i))^{2}$ is asymptotically $T/n$ for doubly-stochastic $W$. We get an explicit upper bound from Thm. 15:
$(\Delta_{T}(j\to i))^{2}\;\leq\;\frac{T}{n}+\frac{2}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{2}{\gamma^{4}}.$ (5.4)
To guarantee $(\varepsilon,\delta)$-DP, we set the Gaussian noise as a function of $\varepsilon$ and $\delta$ via the standard condition (Dwork and Roth, 2014)
$\sigma_{\mathrm{DP}}\;=\;\frac{\Delta_{T}(j\to i)\,\sqrt{2\log(1.25/\delta)}}{\varepsilon}.$ (5.5)
Combining (5.5) with (5.4) gives the per-step DP noise scale
$\sigma_{\mathrm{DP}}^{2}\;\lesssim\;\frac{\log(1/\delta)}{\varepsilon^{2}}\left(\frac{T}{n}+\frac{1}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{1}{\gamma^{4}}\right).$ (5.6)
Gradient Noise for Convergence Bounds of (Koloskova et al., 2020). Because we use full-batch gradients, the only stochasticity is the injected DP noise, and the gradient-noise variances needed for the results of (Koloskova et al., 2020) are given by
$\sigma_{i}^{2}:=\mathbb{E}\|\xi_{i}^{t}\|_{2}^{2}=d\,\sigma_{\mathrm{DP}}^{2},$ (5.7)
which implies the mean noise parameter $\bar{\sigma}^{2}:=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}^{2}=d\,\sigma_{\mathrm{DP}}^{2}.$ Following Koloskova et al. (2020), we also define the heterogeneity measure
$\bar{\zeta}^{2}:=\frac{1}{n}\sum_{i=1}^{n}\|\nabla f_{i}(\theta^{\star})\|_{2}^{2}.$
5.1 Utility Bound for DP Decentralized Learning
We can now directly apply the analysis by Koloskova et al. (2020) by using (5.6) and (5.7). We state the result in a form that the topology appears via $\gamma$ and privacy via $(\varepsilon,\delta)$. More details are given Appendix J.
Theorem 17 (Utility of Full-Batch DP-GD Training with Gossip Averaging).
Assume $f_{i}$ are $L$-smooth and $\mu$-strongly convex, and $W$ is symmetric, doubly-stochastic with absolute spectral gap $\gamma$. Run (5.1) for $T$ rounds with constant stepsize $\eta\simeq 1/L$ and Gaussian noise calibrated so that any node $j$ is $(\varepsilon,\delta)$-DP from the point of view of any node $i$, using (5.5) with $\Delta_{T}(j\to i)$ and the finite bound (5.4). Then the time-averaged excess risk of the average iterate satisfies
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{\theta}^{t})-f^{\star}\right]\;\leq$ $\widetilde{O}\!\bigg(\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n^{2}\,\mu}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{2}\sqrt{T}}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{4}T}$
$\quad\;+\;\frac{L\,\bar{\zeta}^{2}}{\mu^{2}\,\gamma^{2}\,T^{2}}\;+\;\frac{L\|\bar{\theta}^{0}-\theta^{\star}\|_{2}^{2}}{\gamma}\exp\!\Big(-c\,\gamma\frac{\mu}{L}T\Big)\bigg),$
[…]
Proof on a High Level. Theorem 17 is obtained by: (i) viewing full-batch DP gradient descent with gossip averaging as decentralized SGD with gradient noise variance $\bar{\sigma}^{2}=d\sigma_{\mathrm{DP}}^{2}$, (ii) applying the strongly-convex convergence bound by Koloskova et al. (2020) with network parameter $p$ and substituting $p=\Theta(\gamma)$, and (iii) determining $\sigma_{\mathrm{DP}}$ via our single-node-view sensitivity given in Section 4.
Interpretation and Comparison to Centralized DP. The leading term in (17) is a privacy-induced error term $\Theta\!\left(\frac{d\log(1/\delta)}{\varepsilon^{2}n^{2}\mu}\right),$ which matches the scaling one has in the centralized setting (Bassily et al., 2014, 2019, 2020) (see also Sec. 4.2, Ponomareva et al., 2023). The remaining terms are transient and topology-dependent via the absolute spectral gap $\gamma$, vanishing as $T$ grows. Thus, the topology affects how quickly one reaches the privacy-utility baseline of the centralized setting.
6 Conclusions
We presented a framework for analyzing privacy guarantees for DP gossip averaging that accounts for all injected noise, and showed that private gossip averaging gives asymptotically privacy guarantees comparable to central aggregation in case of a single-node view which further leads to utility results comparable to those of centralized private convex learning. We also analyzed the case of multiple-node view and SecLDP, the scenario where the noise injected by the viewing nodes does not contribute to the privacy guarantees. Future work includes analyzing other decentralized methods with similar linear dynamics to assess their long-term privacy guarantees and privacy-utility trade-offs.
Appendix A Adaptive Compositions and Connection to Matrix-Mechanism Accounting
The recent work by Bellet et al. (2025) shows that the linear systems perspective given in Theorem 12 can be equivalently analyzed via a so-called matrix mechanisms (Kairouz et al., 2021b; Pillutla et al., 2025). This analogy allows also extending the non-adaptive privacy analysis directly to adaptive privacy analysis. We next shortly discuss this connection.
Different Threat Scenarios. The matrix mechanism formulation by Bellet et al. (2025) allows analyzing private decentralized optimization methods that have a linear structure. Similarly to our linear systems perspective, it also allows analyzing several existing threat models from the literature, namely:
  • The LDP approach where the adversary sees all messages (Bellet et al., 2018b). This corresponds to the setting of Theorem 12 with $S=I_{n}$.
  • The pair-wise Nerwork Differential Privacy (PNDP) where the adversary is a node and see the messages it sends or receives (Cyffers et al., 2022). This corresponds to Theorem 12 with a selector matrix of a finite collection of nodes.
  • SecLDP (Allouah et al., 2024), where the adversary can remove the randomization it contributes to the computations, thus weakening the privacy guarantees. This corresponds to deletion of certain columns in each block of the matrix $H_{T}$ given in Theorem 12, see Section 3.3.
Privacy Analysis of Matrix Mechanisms. The matrix-mechanism analysis by Bellet et al. (2025) provides a unified viewpoint for characterizing privacy leakage in decentralized learning across different trust settings. The key observation is that the view of any adversarial coalition can be expressed as a linear measurement of a stacked vector of (possibly adaptive and time-varying) messages $G$, corrupted by a structured Gaussian noise term $Z$.
In the matrix mechanism formulation, the $d$-dimensional user-wise messages are stacked into a global matrix $G\in\mathbb{R}^{nT\times d}$ of the form
$G=\begin{bmatrix}G_{1}\\ \vdots\\ G_{T}\end{bmatrix}$,
where each block $G_{t}\in\mathbb{R}^{n\times d}$ corresponds to round $t$ and is given by
$G_{t}=\begin{bmatrix}g_{t}^{1}\\ \vdots\\ g_{t}^{n}\end{bmatrix}$,
with $g_{t}^{i}$ representing the message of node $i$ at iteration $t$. Thus, differences between neighboring datasets appear as differences along the rows $(g_{t}^{i})_{t=1}^{T}$ corresponding to a single user.
The privacy analyses of different private optimization methods under varying threat scenarios are then analyzed via the following results.
Theorem A.1 (Bellet et al., 2025, Thm. 5).
For each of the three trust models, LDP, PNDP and SecLDP, and for every linear decentralized learning algorithm, there exist matrices $A$, $B$ and $C$ such that the adversarial view can be written as
$\mathcal{O}_{\mathcal{A}}=AG+BZ,$ (A.1)
with $A=BC$.
To quantify privacy, one needs to bound how much $G$ may vary under the adjacency relation while remaining observable through this linear mapping.
Definition A.2 (Sensitivity under participation).
Let $B$ and $C$ be two matrices, and let $\Pi$ be a participation scheme for $G$. The sensitivity of $C$ with respect to $B$ under $\Pi$ is defined as
$\mathrm{sens}_{\Pi}(C;B):=\max_{G\simeq_{\Pi}G^{\prime}}\,\|\,C(G-G^{\prime})\,\|_{B^{+}B},$
where $\|v\|_{B^{+}B}:=\sqrt{v^{\top}(B^{+}B)v}$ denotes the seminorm induced by $B^{+}B$, and $G\simeq_{\Pi}G^{\prime}$ indicates that $G$ and $G^{\prime}$ differ only along entries that belong to the same individual according to $\Pi$.
A participation scheme $\Pi$ specifies which coordinates of the stacked update matrix $G$ may change under a neighboring-dataset relation. In other words, $\Pi$ encodes which entries of $G$ correspond to contributions of the same individual across time, ensuring that sensitivity is computed only over data differences permitted by the privacy definition.
Theorem A.3 (Bellet et al., 2025, Thm. 7).
Let $\mathcal{O}_{\mathcal{A}}=AG+BZ$ be the attacker knowledge of a trust model on a linear decentralized learning algorithm, and denote by $\mathcal{M}(G)$ the corresponding mechanism. Let $\Pi$ be a participation scheme for $G$, and suppose $Z\sim\mathcal{N}(0,\nu^{2}I)$. Assume that $A$ is a column-echelon matrix and that $A=BC$ for some $C$. Then, if
$\nu=\sigma\,\mathrm{sens}_{\Pi}(C;B)\qquad\text{with}\qquad\mathrm{sens}_{\Pi}(C;B)\;\leq\;\max_{\pi\in\Pi}\;\sum_{s,t\in\pi}\bigl|\bigl(C^{\top}B^{+}BC\bigr)_{s,t}\bigr|,$
the mechanism $\mathcal{M}$ satisfies $\tfrac{1}{\sigma}$-Gaussian Differential Privacy, even when $G$ is chosen adaptively.
Adaptive Analysis. Notably, Bellet et al. (2025) extend the existing adaptive analysis of matrix mechanisms (Denisov et al., 2022) which apply to square coefficient matrices $A$ to the case where $A$ has less rows than columns. The crucial implication is that once the representation
$\mathcal{O}_{\mathcal{A}}=AG+BZ$
is identified, privacy guarantees for adaptive compositions are the same as in the case of non-adaptive compositions and directly follow from a single sensitivity bound on $(B,C)$ under the participation scheme $\Pi$.
Matrix Mechanism Formulation of Gossip Averaging. We focus on the canonical case of gossip averaging of which DP analysis is equivalent to that of the Muffliato-SGD with a single Muffliato averaging step. As shown in (Appendix A.4 Bellet et al., 2025), in this case the matrix mechanism representation is determined by the matrices $A=B=H_{T}$ and $C=I$, where $H_{T}$ is the matrix given in Thm. 12.
Equivalence of Sensitivity Bounds of Thm. 12 and of (Bellet et al., 2025).We also see from Thm. A.3 that the sensitivity upper bound equals the upper bound given in our Thm. 12 in case the differing node participates in every round, since if $H_{T}=U\Sigma V^{T}$ is the compact SVD of $H_{T}$, $B=H_{T}$ and $C=I$, then
$\left|C^{\top}B^{+}BC\right|=\left|VV^{\top}\right|$
and the summation in the upper bound of Thm. A.3 equals the upper bound given in our Thm. 12 in case the differing node participates in every round.
Appendix B Proof of Lemma 6
Proof.
Recall, the projected Gaussian mechanism is give as
$\mathcal{M}(D)=f(D)+AZ,$
where $f:\mathcal{D}\rightarrow\mathbb{R}^{m}$, $A\in\mathbb{R}^{m\times n}$ and $Z\sim\mathcal{N}(0,\sigma^{2}I_{n})$ for some $\sigma>0$.
Denote the compact SVD of $A$ as $A=U_{r}\Sigma_{r}V_{r}^{\top}$, where $r$ denotes the rank of $A$. Since the columns of $U_{r}$ give a basis for the subspace $\mathrm{Range(A)}$ and $U_{r}$ has orthonormal columns, $U_{r}U_{r}^{\top}$ gives a projector onto $\mathrm{Range(A)}$. Since $f(D)-f(D^{\prime})\in\mathrm{Range(A)}$, it holds that
$f(D)-f(D^{\prime})=U_{r}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big).$ (B.1)
By using the compact SVD of $A$, we have that for all $\alpha\geq 0$:
$H_{\alpha}\left(\mathcal{M}(D)||\mathcal{M}(D^{\prime})\right)$ $=H_{\alpha}\left(f(D)+AZ\,||\,f(D^{\prime})+AZ\right)$
$=H_{\alpha}\left(f(D)-f(D^{\prime})+AZ\,||\,AZ\right)$
$=H_{\alpha}\left(f(D)-f(D^{\prime})+U_{r}\Sigma_{r}V_{r}^{\top}Z\,||\,U_{r}\Sigma_{r}V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(U_{r}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)+U_{r}\Sigma_{r}V_{r}^{\top}Z\,||\,U_{r}\Sigma_{r}V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)+\Sigma_{r}V_{r}^{\top}Z\,||\,\Sigma_{r}V_{r}^{\top}Z\right)$
$=H_{\alpha}\left(\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)+V_{r}^{\top}Z\,||\,V_{r}^{\top}Z\right)$
[…]
where $\widetilde{Z}\sim\mathcal{N}(0,\sigma^{2}I_{r})$ and where before the third last equality we have carried out multiplication from the left by $U_{r}^{\top}$ and before the second last equality we have carried out multiplication from the left by $\Sigma_{r}^{-1}$. The last step follows from the fact that $V_{r}$ has orthonormal columns. We see that for all $\varepsilon\in\mathbb{R}$,
[…]
where the right-hand side gives the $\delta(\varepsilon)$-privacy profile for the Gaussian mechanism with sensitivity $\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ and noise scale $\sigma$. Further, since $V_{r}$ has orthonormal columns,
$\left\lVert\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$ $=\left\lVert V_{r}\Sigma_{r}^{-1}U_{r}^{\top}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2}$
$=\left\lVert A^{+}\big(f(D)-f(D^{\prime})\big)\right\rVert_{2},$
and the claim follows. ∎
Appendix C Proof of Lemma 9
Proof.
Since $W$ is symmetric, $W=Q\Lambda Q^{\top}$ with $Q$ orthogonal and $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})$ real. Doubly stochastic implies $\lambda_{1}=1$ with eigenvector $\mathbf{1}$. Primitivity implies $|\lambda_{k}|<1$ for all $k\geq 2$. Let $P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ be the orthogonal projector onto $\mathrm{span}\{\mathbf{1}\}$ and set R:=WPR:=W-P. Then $RP=PR=0$ and $\|R^{t}\|_{2}=\max_{k\geq 2}|\lambda_{k}|^{t}$. With $\rho:=\max_{k\geq 2}|\lambda_{k}|\in(0,1)$ the claim follows. ∎
Appendix D Proof of Lemma 10
Proof.
From the translational invariance of the hockey-stick divergence, we have
$H_{\alpha}\left(\mathrm{View}_{\mathcal{M}(D)}(\{i\})\,||\,\mathrm{View}_{\mathcal{M}(D^{\prime})}(\{i\}))\right)$ $=H_{\alpha}\left(H_{T}\widehat{x}_{T}+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widehat{x}_{T}^{\prime}+H_{T}\widetilde{u}_{T}\right)$
$=H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right)$
which gives the claim.
Appendix E Proof of Thm. 12
Proof.
The claim is obtained by applying Lemma 6 (DP guarantees of a projected Gaussian mechanism via Moore–Penrose pseudoinverse) to the characterization given by Lemma 10. For completeness, we apply the pseudoinverse $H_{T}^{+}$ step by step. By Lemma 10, the privacy guarantees are obtained via the hockey-stick divergence $H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right)$, $\alpha\geq 0$. By the compact SVD $H_{T}=U_{r}\Sigma_{r}V_{r}^{\top}$, where $r$ denotes the rank of $H_{T}$, we have that for all $\alpha\geq 0$,
$H_{\alpha}\left(H_{T}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+H_{T}\widetilde{u}_{T}\,||\,H_{T}\widetilde{u}_{T}\right)$
$=H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
$=H_{\alpha}\left(U_{r}\Sigma_{r}V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\,||\,U_{r}\Sigma_{r}V_{r}^{\top}\widetilde{u}_{T}\right)$
$=H_{\alpha}\left(V_{r}^{\top}\big(\widehat{x}_{T}-\widehat{x}_{T}^{\prime}\big)+V_{r}^{\top}\widetilde{u}_{T}\,||\,V_{r}^{\top}\widetilde{u}_{T}\right)$
[…]
where
$c=\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}$
and $\widehat{u}\sim\mathcal{N}(0,\sigma^{2}I_{T})$. ∎
For an upper bound of the sensitivity, we need to find the maximum of $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ when $\Delta x$ is in the convex and compact domain
$\mathcal{D}_{j}:=\{\alpha\otimes e_{j}:\alpha\in[-1,1]^{T}\}$
We see that maximizing $\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}$ is equivalent to finding the maximum of the quadratic form $(\Delta x)^{\top}P\Delta x$ in $\mathcal{D}_{j}$ where the projector matrix $P=V_{r}V_{r}^{\top}$ is positive semi-definite and a projection on to the row space of $H_{T}$. Thus, the optimum is found from somewhere at the boundary of $\mathcal{D}_{j}$, i.e., it holds that $\Delta x=\alpha\otimes e_{j}$, where $\alpha=\{-1,1\}^{T}$.
For the inequality, we use the trivial inequality
$\left\lVert V_{r}^{\top}\Delta x\right\rVert_{2}^{2}=(\Delta x)^{\top}V_{r}V_{r}^{\top}(\Delta x)\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|V_{r}V_{r}^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$
where $\left|\cdot\right|$ denotes the element-wise abolute value of a matrix. ∎
Appendix F Numerical Illustration of the Spectral Estimate of Sec. 3.4
We next consider a numerical illustration of the upper bound for the sensitivity
$\Delta^{T}_{j\rightarrow\mathcal{A}}=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2},$
where $H_{T}$ is the system matrix given in Lemma 10. We compare the upper bound based on the spectral estimate $\lambda_{\max}(M_{T})$ to a lower bound for the sensitivity $\Delta^{T}_{j\rightarrow\mathcal{A}}$ we obtain with the choice $c=\mathbf{1}=\begin{bmatrix}1&\ldots&1\end{bmatrix}^{\top}$. We compare the bounds in the same two cases considered in Section 4.1: we first consider a randomly drawn synthetic Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$, and second, a network generated using a preferential-attachment model (Barabási and Albert, 1999) with $n=100$ nodes, initialized from a fully connected core of 5 nodes, where each new node forms 3 links to existing nodes, with connection probabilities proportional to current node degrees. And in both cases, we take the gossip matrix $W$ to be a doubly stochastic matrix obtained from so called Metropolis–Hastings averaging of a given symmetric adjacency matrix.
As illustrated by Figure 2 the upper bound is in both cases less than 10 percent higher than the actual tight upper bound.
Figure 2: The upper bound for the squared and scaled sensitivity $(\Delta^{T}(j\to i))^{2}/T$ as a function of $T$ for a randomly drawn Erdős–Rényi graph $G(n,p)$ with $n=100$ and $p=0.15$ and b) a randomly drawn doubly-stochastic gossip matrix of a preferential-attachment model with $n=100$ nodes.
Appendix G Proof of Theorem 13
We next give proof of Theorem 13. In the proof, Lemma G.2 (Section G.2) contains a crucial step: it shows that the transient component of the sensitivity can be bounded independent of $T$.
G.1 Proof of Theorem 13
Theorem G.1.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric, doubly stochastic, and primitive. Let $S=e_{i}^{\top}$ be the single-node observation matrix. Then for any $j\in\{1,\dots,n\}$,
$\lim_{T\to\infty}\frac{\Delta_{T}(j\to i)^{2}}{T}=\frac{1}{n},$
where
$\Delta_{T}(j\to i)^{2}:=\max_{c\in\{\pm 1\}^{T}}\left\|H_{T}^{+}H_{T}(c\otimes e_{j})\right\|_{2}^{2}.$
Proof.
We start by stating the sensitivity formula as a quadratic form. For $c\in\{\pm 1\}^{T}$ define
$v(c):=c\otimes e_{j},\qquad b(c):=H_{T}v(c)\in\mathbb{R}^{T},$
and set
$K_{T}:=H_{T}H_{T}^{\top}.$
Since $H_{T}^{+}=H_{T}^{\top}(H_{T}H_{T}^{\top})^{-1}$ and $H_{T}$ has full row rank, $K_{T}\succ 0$ and
$\|H_{T}^{+}H_{T}v(c)\|_{2}^{2}=b(c)^{\top}K_{T}^{-1}b(c),$
hence
$\Delta_{T}(j\to i)^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}K_{T}^{-1}b(c).$ (G.1)
Denoting $a_{m}:=(W^{m})_{ij}$ for $m\in\mathbb{N}$, we have for $t=1,\dots,T$,
$b_{t}(c)=\sum_{k=1}^{t}a_{t-k}c_{k}.$ (G.2)
We start by decomposing $K_{T}$. By Lemma 9, we can write
$W=P+R,\qquad P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top},\qquad\mathbf{1}^{\top}R=0,\qquad RP=PR=0,$ (G.3)
and $\|R^{t}\|_{2}\leq\rho^{t}$ for $\rho:=\max_{k\geq 2}|\lambda_{k}(W)|$. Using (G.3) together with the block-Toeplitz structure of $H_{T}$, we obtain the decomposition
$K_{T}=\frac{1}{n}J+G_{T},$ (G.4)
where $G_{T}\succeq 0$ and $J\succ 0$ satisfies $J=LL^{\top}$, with $L$ the $T\times T$ lower-triangular all-ones matrix:
$L=\begin{bmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 1&1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&1&1&\cdots&1\end{bmatrix}\in\mathbb{R}^{T\times T}.$
Equivalently, $J_{t,s}=\min(t,s)$. Therefore,
$K_{T}\succeq\frac{1}{n}J$
which further implies that
$K_{T}^{-1}\preceq nJ^{-1}.$ (G.5)
Let $D$ be the backward difference matrix on $\mathbb{R}^{T}$,
$(Dx)_{1}=x_{1},\qquad(Dx)_{t}=x_{t}-x_{t-1}\ \ (t\geq 2),$
i.e.,
$D=\begin{bmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&0&\cdots&-1&1\end{bmatrix}\in\mathbb{R}^{T\times T}.$
Since $L^{-1}=D$ and $J^{-1}=D^{\top}D$, for all $x\in\mathbb{R}^{T}$,
$x^{\top}J^{-1}x=\|Dx\|_{2}^{2}.$ (G.6)
We next split $b(c)$ into stationary and transient parts. Define the stationary and transient terms as
$a_{m}^{(P)}:=(P)_{ij}=\frac{1}{n}\quad(m\geq 1),\qquad a_{m}^{(R)}:=(R^{m})_{ij}\quad(m\geq 1),$ (G.7)
and note $a_{0}=(W^{0})_{ij}=\mathbf{1}\{i=j\}$. Accordingly, write
$b(c)=b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c),$ (G.8)
where $b^{(0)}$ denotes the (possible) $a_{0}$ contribution,
$b_{t}^{(P)}(c):=\sum_{k=1}^{t-1}\frac{1}{n}\,c_{k}=\frac{1}{n}\sum_{k=1}^{t-1}c_{k},$ (G.9)
and
$b_{t}^{(R)}(c):=\sum_{k=1}^{t-1}(R^{t-k})_{ij}c_{k}=\sum_{m=1}^{t-1}(R^{m})_{ij}\,c_{t-m}.$ (G.10)
We next derive an upper bound for the quadratic form $b(c)^{\top}K_{T}^{-1}b(c)$. For any $c$,
$b(c)^{\top}K_{T}^{-1}b(c)$ $=\|b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c)\|_{K_{T}^{-1}}^{2}$
$\leq\Big(\|b^{(0)}(c)\|_{K_{T}^{-1}}+\|b^{(P)}(c)\|_{K_{T}^{-1}}+\|b^{(R)}(c)\|_{K_{T}^{-1}}\Big)^{2}.$
The term $b^{(0)}(c)$ contributes at most an $O(1)$ additive constant, so we ignore it in the asymptotics. By (G.5) and (G.6),
$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}\leq n\,\|Db^{(P)}(c)\|_{2}^{2}.$
From (G.9), $(Db^{(P)}(c))_{t}=\frac{1}{n}c_{t-1}$ for $t\geq 2$ and $(Db^{(P)}(c))_{1}=0$, hence $\|Db^{(P)}(c)\|_{2}^{2}=(T-1)/n^{2}$ and therefore
$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}\leq\frac{T-1}{n}.$ (G.11)
By Lemma G.2 given in Section G.2, $\|b^{(R)}(c)\|_{K_{T}^{-1}}^{2}\leq C_{\mathrm{mix}}(W)$. Thus, for all $c$,
$b(c)^{\top}K_{T}^{-1}b(c)\leq\left(\sqrt{\frac{T-1}{n}}+\sqrt{C_{\mathrm{mix}}(W)}+O(1)\right)^{2}=\frac{T}{n}+O(\sqrt{T}).$
Taking the maximum over $c$ gives
$\limsup_{T\to\infty}\frac{\Delta_{T}(j\to i)^{2}}{T}\leq\frac{1}{n}.$ (G.12)
We derive a lower bound using $c=\mathbf{1}$, i.e., $c$ being all ones. Then
$(Db)_{t}=b_{t}-b_{t-1}=(W^{t-1})_{ij}=\frac{1}{n}+r_{t-1}.$
Hence
$\|Db\|_{2}^{2}=\sum_{t=1}^{T}\left(\frac{1}{n}+r_{t-1}\right)^{2}=\frac{T}{n^{2}}+\frac{2}{n}\sum_{t=1}^{T}r_{t-1}+\sum_{t=1}^{T}r_{t-1}^{2}.$
Using $|r_{t-1}|\leq\rho^{t-1}$ and summing geometric series, we have
$\left|\sum_{t=1}^{T}r_{t-1}\right|\leq\sum_{t=1}^{\infty}\rho^{t-1}=\frac{1}{1-\rho}$
and
$\sum_{t=1}^{T}r_{t-1}^{2}\leq\sum_{t=1}^{\infty}\rho^{2(t-1)}=\frac{1}{1-\rho^{2}}.$
Therefore,
$\|Db\|_{2}^{2}\geq\frac{T}{n^{2}}-\frac{2}{n(1-\rho)}-\frac{1}{1-\rho^{2}}.$ (G.13)
Next, recall that
$K_{T}=H_{T}H_{T}^{\top}=\frac{1}{n}J+G_{T},$
where $J=LL^{\top}$ with $J_{t,s}=\min(t,s)$, and $G_{T}\succeq 0$, and $J^{-1}=D^{\top}D$, where $D$ is the backward difference operator.
Since $K_{T}$ is positive definite, we use the variational representation
$b^{\top}K_{T}^{-1}b=\max_{x\in\mathbb{R}^{T}}\left(2b^{\top}x-x^{\top}K_{T}x\right).$
We lower bound this expression using
$x_{0}:=nJ^{-1}b=nD^{\top}Db.$
Then
$b^{\top}K_{T}^{-1}b$ $\geq 2b^{\top}x_{0}-x_{0}^{\top}K_{T}x_{0}$
$=2n\,b^{\top}J^{-1}b-x_{0}^{\top}\left(\tfrac{1}{n}J+G_{T}\right)x_{0}$
$=n\,b^{\top}J^{-1}b-x_{0}^{\top}G_{T}x_{0}$
$=n\|Db\|_{2}^{2}-x_{0}^{\top}G_{T}x_{0}.$
Thus, the required task is to bound the term $x_{0}^{\top}G_{T}x_{0}$. Write
$W=\frac{1}{n}11^{\top}+R$
with $1^{\top}R=0$ and $\|R^{t}\|_{2}\leq\rho^{t}$. This gives a decomposition
$H_{T}=H_{T}^{1}+H_{T}^{R},$
where $H_{T}^{1}$ is generated by $\frac{1}{n}11^{\top}$ and $H_{T}^{R}$ by $R$. Consequently,
$G_{T}=H_{T}^{R}(H_{T}^{R})^{\top}.$
Since $\|R^{t}\|_{2}\leq\rho^{t}$, a geometric-series argument gives (the 2-norm if defined by the first block column of the block lower-triangular matrix $H_{T}^{R}$)
$\|H_{T}^{R}\|_{2}\leq\frac{1}{1-\rho}.$
Recall that
$x_{0}=nD^{\top}Db,\qquad(Db)_{t}=\frac{1}{n}+r_{t-1},\qquad|r_{t}|\leq\rho^{t}.$
Let $y:=Db$. We then have that
$(D^{\top}y)_{t}=y_{t}-y_{t+1}\quad(1\leq t\leq T-1),\qquad(D^{\top}y)_{T}=y_{T}.$
Hence, for $1\leq t\leq T-1$,
$(x_{0})_{t}=n(y_{t}-y_{t+1})=n(r_{t-1}-r_{t}),$
and
$(x_{0})_{T}=ny_{T}=1+nr_{T-1}.$
Using $|r_{t-1}-r_{t}|\leq|r_{t-1}|+|r_{t}|\leq(1+\rho)\rho^{t-1}$, we obtain
$\sum_{t=1}^{T-1}(x_{0})_{t}^{2}\leq n^{2}(1+\rho)^{2}\sum_{t=1}^{\infty}\rho^{2(t-1)}=n^{2}\frac{(1+\rho)^{2}}{1-\rho^{2}}=n^{2}\frac{1+\rho}{1-\rho}.$
Moreover,
$(x_{0})_{T}^{2}=(1+nr_{T-1})^{2}\leq(1+n)^{2}.$
Combining these gives
$\|x_{0}\|_{2}^{2}\leq(1+n)^{2}+n^{2}\frac{1+\rho}{1-\rho},$
and furthermore
$x_{0}^{\top}G_{T}x_{0}\leq\frac{(1+n)^{2}}{(1-\rho)^{2}}+n^{2}\frac{1+\rho}{(1-\rho)^{3}}.$ (G.14)
Combining the bounds of Eq. (G.13) and (G.14) gives
$b^{\top}K_{T}^{-1}b\geq\frac{T}{n}-\frac{2}{n(1-\rho)}-\frac{1}{1-\rho^{2}}-\frac{(1+n)^{2}}{(1-\rho)^{2}}-n^{2}\frac{1+\rho}{(1-\rho)^{3}}.$
Since the particular choice of $c=\mathbf{1}$ gives a lower bound, we see that
$\limsup_{T\to\infty}\frac{\Delta_{T}(j\to i)^{2}}{T}\geq\frac{1}{n}.$
This together with (G.12) completes the proof. ∎
G.2 Bound for the Transient Term
We next prove a technical lemma required for the proof of Thm. 13. This lemma shows that the transient terms stay bounded independent of $T$.
Lemma G.2 (Uniform bound for the transient component).
Let $c\in\{-1,1\}^{T}$ and let $b^{(R)}(c)$ and $K_{T}$ be defined as above in the proof of Thm. 13. There exists a constant $C_{\mathrm{mix}}(W)<\infty$ depending only on $W$, such that for all $T\geq 1$ and all $c\in\{\pm 1\}^{T}$,
$b^{(R)}(c)^{\top}K_{T}^{-1}b^{(R)}(c)\leq C_{\mathrm{mix}}(W),$
where $C_{\mathrm{mix}}(W)=\frac{\pi^{2}}{6}(1-\rho)^{-4}$ and $\rho:=\max_{k\geq 2}|\lambda_{k}|$.
Proof.
For the proof we repeatedly use the Schur complement (Gallier, 2010). For $t=1,\dots,T$, let $K_{t}$ be the leading $t\times t$ principal submatrix of $K_{T}$ and write
$K_{t}=\begin{pmatrix}K_{t-1}&k_{t}\\ k_{t}^{\top}&\kappa_{t}\end{pmatrix},\qquad s_{t}:=\kappa_{t}-k_{t}^{\top}K_{t-1}^{-1}k_{t}>0.$
For any $x\in\mathbb{R}^{T}$ and $t\in[T]$, define the residual
$r_{t}(x):=x_{t}-k_{t}^{\top}K_{t-1}^{-1}x_{1:t-1}.$
A standard block inverse identity with Schur complement gives
$x^{\top}K_{T}^{-1}x=\sum_{t=1}^{T}\frac{r_{t}(x)^{2}}{s_{t}}.$ (G.15)
We next lower bound $s_{t}$ uniformly. Let $h_{t}^{\top}$ be the $t$th row of $H_{T}$. Then $K_{T}$ is the Gram matrix of $\{h_{t}\}_{t=1}^{T}$, and $s_{t}$ equals the squared distance from $h_{t}$ to $\mathrm{span}\{h_{1},\dots,h_{t-1}\}$. Since $H_{T}$ is block lower-triangular with diagonal block $S=e_{i}^{\top}$, each diagonal component is orthogonal to $\mathrm{span}\{h_{1},\dots,h_{t-1}\}$ and we clearly have that
$s_{t}\geq\|S\|_{2}^{2}=1.$ (G.16)
Combining (G.15)–(G.16) gives
$x^{\top}K_{T}^{-1}x\leq\sum_{t=1}^{T}r_{t}(x)^{2}.$ (G.17)
We now bound the residuals for $x=b^{(R)}(c)$. Since $k_{t}^{\top}K_{t-1}^{-1}$ is the minimizer of $\alpha^{\top}\mapsto(x_{t}-\alpha^{\top}x_{1:t-1})^{2}$, for any $\widehat{x}_{t}=\alpha^{\top}x_{1:t-1}$,
$r_{t}(x)^{2}\leq(x_{t}-\widehat{x}_{t})^{2}.$ (G.18)
Fix $t\geq 2$ and choose a truncation length
$m_{t}:=\left\lceil\frac{\log t}{\log(1/\rho)}\right\rceil\quad\Rightarrow\quad\rho^{m_{t}}\leq\frac{1}{t}.$
Define a predictor $\widehat{b}^{(R)}_{t}$ as
$\widehat{b}^{(R)}_{t}:=\sum_{q=1}^{m_{t}}\beta_{q}\,b^{(R)}_{t-q}(c),\qquad\beta_{q}:=(W^{q})_{ii}.$ (G.19)
We show that the error $e_{t}:=b^{(R)}_{t}-\widehat{b}^{(R)}_{t}$ is $O\!\big(\rho^{m_{t}}/(1-\rho)\big)$. Using (G.10) we may write
$b^{(R)}_{t}(c)=\sum_{m=1}^{t-1}r_{m}\,c_{t-m},\qquad r_{m}:=(R^{m})_{ij}.$
Similarly,
$b^{(R)}_{t-q}(c)=\sum_{m=1}^{t-q-1}r_{m}\,c_{t-q-m}=\sum_{\ell=q+1}^{t-1}r_{\ell-q}\,c_{t-\ell},$
and therefore
$\widehat{b}^{(R)}_{t}=\sum_{q=1}^{m_{t}}\beta_{q}\sum_{\ell=q+1}^{t-1}r_{\ell-q}\,c_{t-\ell}=\sum_{\ell=2}^{t-1}\left(\sum_{q=1}^{\min\{m_{t},\ell-1\}}\beta_{q}r_{\ell-q}\right)c_{t-\ell}.$
Hence
$e_{t}=\sum_{\ell=1}^{t-1}\left(r_{\ell}-\sum_{q=1}^{\min\{m_{t},\ell-1\}}\beta_{q}r_{\ell-q}\right)c_{t-\ell}.$ (G.20)
Now use the identity $W^{q}R^{\ell-q}=R^{\ell}$ for all $1\leq q\leq\ell-1$, which follows from $W=P+R$ and $PR=RP=0$. Taking the $(i,j)$ entry,
$r_{\ell}=(R^{\ell})_{ij}=\sum_{u=1}^{n}(W^{q})_{iu}\,(R^{\ell-q})_{uj}.$ (G.21)
Recalling $\beta_{q}=(W^{q})_{ii}$ gives
$r_{\ell}-\beta_{q}r_{\ell-q}=\sum_{u\neq i}(W^{q})_{iu}\,(R^{\ell-q})_{uj}.$ (G.22)
Summing (G.22) over $q=1,\dots,\min\{m_{t},\ell-1\}$ and plugging into (G.20), we see that for all $\ell\geq m_{t}+1$ the coefficient of $c_{t-\ell}$ is a sum of terms $(W^{q})_{iu}(R^{\ell-q})_{uj}$ with $u\neq i$ and $1\leq q\leq m_{t}$. Using row-stochasticity ($\sum_{u\neq i}(W^{q})_{iu}\leq 1$) and $|(R^{k})_{uj}|\leq\|R^{k}\|_{2}\leq\rho^{k}$, we obtain
$\left|r_{\ell}-\sum_{q=1}^{m_{t}}\beta_{q}r_{\ell-q}\right|\leq\sum_{q=1}^{m_{t}}\sum_{u\neq i}(W^{q})_{iu}\,|(R^{\ell-q})_{uj}|\leq\sum_{q=1}^{m_{t}}\rho^{\ell-q}\leq\frac{\rho^{\ell-m_{t}}}{1-\rho}.$
For $\ell\leq m_{t}$ the bracket in (G.20) is uniformly bounded by
$|r_{\ell}|+\sum_{q\leq\ell-1}|\beta_{q}||r_{\ell-q}|\leq\rho^{\ell}+\sum_{q\leq\ell-1}\rho^{\ell-q}\leq 1/(1-\rho).$
Combining these bounds and using $|c_{t-\ell}|=1$ gives
$|e_{t}|\leq\sum_{\ell=m_{t}+1}^{t-1}\frac{\rho^{\ell-m_{t}}}{1-\rho}\leq\frac{1}{(1-\rho)^{2}}\rho^{m_{t}}\leq\frac{1}{(1-\rho)^{2}}\frac{1}{t}.$ (G.23)
Finally, by (G.18) and (G.23),
$r_{t}(b^{(R)}(c))^{2}\leq e_{t}^{2}\leq\frac{1}{(1-\rho)^{4}}\frac{1}{t^{2}}.$ (G.24)
Using (G.17), summing over $t$ and using the bound (G.24) gives
$b^{(R)}(c)^{\top}K_{T}^{-1}b^{(R)}(c)\leq\sum_{t=1}^{T}r_{t}(b^{(R)}(c))^{2}\leq\frac{1}{(1-\rho)^{4}}\sum_{t=1}^{\infty}\frac{1}{t^{2}}\leq\frac{\pi^{2}/6}{(1-\rho)^{4}}.$
This proves the claim. ∎
Appendix H Extension to Multiple Observed Nodes
We next give proof of Theorem 15. The proof follows exactly the same structure as the proof of Theorem 13: we express the sensitivity as a quadratic form, decompose the dynamics into a stationary component and a transient component, show that the stationary part gives the term linear in $T$, and prove via an auxiliary technical lemma that transient contributions are uniformly bounded in $T$.
H.1 Proof of Theorem 15
Proof.
We start by stating the sensitivity formula as a quadratic form. For $c\in\{\pm 1\}^{T}$ define
$v(c):=c\otimes e_{j},\qquad b(c):=H_{T}v(c)\in\mathbb{R}^{mT},$
and set
$K_{T}:=H_{T}H_{T}^{\top}.$
Since $H_{T}^{+}=H_{T}^{\top}(H_{T}H_{T}^{\top})^{-1}$ and $H_{T}$ has full row rank, $K_{T}\succ 0$ and
$\|H_{T}^{+}H_{T}v(c)\|_{2}^{2}=b(c)^{\top}K_{T}^{-1}b(c),$
hence
$\Delta_{T}(j\to\mathcal{A})^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}K_{T}^{-1}b(c).$ (H.1)
Denote
$a_{\ell}:=S_{\mathcal{A}}W^{\ell}e_{j}\in\mathbb{R}^{m},\qquad\ell\in\mathbb{N}.$
Then for $t=1,\dots,T$, the $t$th $m$-block of $b(c)$ satisfies the same convolution identity
$b_{t}(c)=\sum_{k=1}^{t}a_{t-k}\,c_{k}.$ (H.2)
We start by decomposing $K_{T}$. By Lemma 9 (Lemma 9 in Appendix C), we can write
$W=P+R,\qquad P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top},\qquad\mathbf{1}^{\top}R=0,\qquad RP=PR=0,$ (H.3)
and $\|R^{t}\|_{2}\leq\rho^{t}$ for $\rho:=\max_{k\geq 2}|\lambda_{k}(W)|\in(0,1)$. Using (H.3) together with the block-Toeplitz structure of $H_{T}$ and the fact that $S_{\mathcal{A}}S_{\mathcal{A}}^{\top}=I_{m}$, one obtains the decomposition
$K_{T}=\frac{1}{n}(J\otimes I_{m})+G_{T},$ (H.4)
where $G_{T}\succeq 0$ and $J\succ 0$ satisfies $J=LL^{\top}$ with $L$ the $T\times T$ lower-triangular all-ones matrix. Equivalently, $J_{t,s}=\min(t,s)$. Therefore,
$K_{T}\succeq\frac{1}{n}(J\otimes I_{m})$
which further implies
$K_{T}^{-1}\preceq n(J^{-1}\otimes I_{m}).$ (H.5)
Let $D$ be the backward difference matrix on $\mathbb{R}^{T}$,
$(Dx)_{1}=x_{1},\qquad(Dx)_{t}=x_{t}-x_{t-1}\ \ (t\geq 2),$
so that $J^{-1}=D^{\top}D$ and for all $x\in\mathbb{R}^{T}$,
$x^{\top}J^{-1}x=\|Dx\|_{2}^{2}.$ (H.6)
We next split $b(c)$ into stationary and transient parts, exactly as in the proof of Thm. 13. Define the stationary and transient terms as
$a_{\ell}^{(P)}:=S_{\mathcal{A}}Pe_{j}=\frac{1}{n}\mathbf{1}_{m}\quad(\ell\geq 1),\qquad a_{\ell}^{(R)}:=S_{\mathcal{A}}R^{\ell}e_{j}\quad(\ell\geq 1),$ (H.7)
and $a_{0}=S_{\mathcal{A}}e_{j}$. Accordingly, write
$b(c)=b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c),$ (H.8)
for $t=1,\dots,T$,
$b_{t}^{(P)}(c):=\sum_{k=1}^{t-1}\frac{1}{n}\,\mathbf{1}_{m}\,c_{k}=\frac{1}{n}\mathbf{1}_{m}\sum_{k=1}^{t-1}c_{k},$ (H.9)
and
$b_{t}^{(R)}(c):=\sum_{m=1}^{t-1}S_{\mathcal{A}}R^{m}e_{j}\,c_{t-m}.$ (H.10)
We next derive an upper bound for the quadratic form $b(c)^{\top}K_{T}^{-1}b(c)$. For any $c$,
$b(c)^{\top}K_{T}^{-1}b(c)$ $=\|b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c)\|_{K_{T}^{-1}}^{2}$
$\leq\Big(\|b^{(0)}(c)\|_{K_{T}^{-1}}+\|b^{(P)}(c)\|_{K_{T}^{-1}}+\|b^{(R)}(c)\|_{K_{T}^{-1}}\Big)^{2}.$
By (H.5) and (H.6),
$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}$ $\leq n\,(b^{(P)}(c))^{\top}(J^{-1}\otimes I_{m})\,b^{(P)}(c)=n\,\|(D\otimes I_{m})b^{(P)}(c)\|_{2}^{2}.$
From (H.9), we have $(Db^{(P)}(c))_{t}=\frac{1}{n}\,\mathbf{1}_{m}\,c_{t-1}$ for $t\geq 2$ and $(Db^{(P)}(c))_{1}=0$, hence
$\|(D\otimes I_{m})b^{(P)}(c)\|_{2}^{2}=\sum_{t=2}^{T}\left\|\frac{1}{n}\,\mathbf{1}_{m}\,c_{t-1}\right\|_{2}^{2}=\sum_{t=2}^{T}\frac{m}{n^{2}}=\frac{m(T-1)}{n^{2}}.$
Therefore
$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}\leq\frac{m(T-1)}{n}.$ (H.11)
By Lemma H.1, $\|b^{(R)}(c)\|_{K_{T}^{-1}}^{2}\leq C_{\mathrm{mix}}(W)$. Thus, for all $c$,
$b(c)^{\top}K_{T}^{-1}b(c)\leq\left(\sqrt{\frac{m(T-1)}{n}}+\sqrt{C_{\mathrm{mix}}(W)}+O(1)\right)^{2}=\frac{m}{n}T+O(\sqrt{T}).$
The constants given in the theorem statement are obtained by combining the $O(1)$ term with the $\sqrt{\frac{m(T-1)}{n}}$-term. ∎
H.2 Bound for the Transient Term
We next prove the required technical lemma which shows that the transient terms stay bounded independent of $T$ also for the multi-node observation.
Lemma H.1.
Let $c\in\{-1,1\}^{T}$ and let $\widetilde{b}^{(R)}(c)$ and $\widetilde{K}_{T}$ be defined as above in the proof of Thm. 15. There exists a constant $C_{\mathrm{mix}}(W)<\infty$ depending only on $W$, such that for all $T\geq 1$ and all $c\in\{\pm 1\}^{T}$,
$\left(\widetilde{b}^{(R)}(c)\right)^{\top}\widetilde{K}_{T}^{-1}\,\widetilde{b}^{(R)}(c)\leq C_{\mathrm{mix}}(W),$ (H.12)
where one may take
$C_{\mathrm{mix}}(W)=\frac{\pi^{2}}{6}\,m\,(1-\rho)^{-4},\qquad\rho:=\max_{k\geq 2}|\lambda_{k}(W)|.$
Proof.
The proof follows the same Schur-complement argument as Lemma G.2, with the only difference that now the vectors live in $\mathbb{R}^{mT}$.
For $t=1,\dots,T$, let $\widetilde{K}_{t}$ be the leading $(mt)\times(mt)$ principal submatrix of $\widetilde{K}_{T}$ and write it in block form
$\widetilde{K}_{t}=\begin{pmatrix}\widetilde{K}_{t-1}&\widetilde{k}_{t}\\ \widetilde{k}_{t}^{\top}&\widetilde{\kappa}_{t}\end{pmatrix},\qquad\widetilde{s}_{t}:=\widetilde{\kappa}_{t}-\widetilde{k}_{t}^{\top}\widetilde{K}_{t-1}^{-1}\widetilde{k}_{t}\succ 0,$
where $\widetilde{\kappa}_{t}\in\mathbb{R}^{m\times m}$ and $\widetilde{s}_{t}\in\mathbb{R}^{m\times m}$. For any $x\in\mathbb{R}^{mT}$ define the residual vector
$\widetilde{r}_{t}(x):=x_{t}-\widetilde{k}_{t}^{\top}\widetilde{K}_{t-1}^{-1}x_{1:t-1}\in\mathbb{R}^{m},$
where $x_{t}\in\mathbb{R}^{m}$ denotes the $t$th block of $x$. A standard block inverse identity with Schur complement gives
$x^{\top}\widetilde{K}_{T}^{-1}x=\sum_{t=1}^{T}\widetilde{r}_{t}(x)^{\top}\widetilde{s}_{t}^{-1}\widetilde{r}_{t}(x).$ (H.13)
We next lower bound $\widetilde{s}_{t}$. Let $\widetilde{h}_{t}^{\top}$ be the $t$th block-row of $\widetilde{H}_{T}$. Then, $\widetilde{K}_{T}$ is the Gram matrix of $\{\widetilde{h}_{t}\}_{t=1}^{T}$. Here $\widetilde{s}_{t}$ admits a geometric interpretation analogous to the scalar case. Since $\widetilde{K}_{T}$ is the Gram matrix of the block-rows $\{\widetilde{h}_{1},\dots,\widetilde{h}_{T}\}$ of $\widetilde{H}_{T}$, the Schur complement
$\widetilde{s}_{t}=\widetilde{\kappa}_{t}-\widetilde{k}_{t}^{\top}\widetilde{K}_{t-1}^{-1}\widetilde{k}_{t}$
is the Gram matrix of the component of $\widetilde{h}_{t}$ orthogonal to $\mathrm{span}\{\widetilde{h}_{1},\dots,\widetilde{h}_{t-1}\}$. Equivalently, for any $z\in\mathbb{R}^{m}$, $z^{\top}\widetilde{s}_{t}z$ equals the squared distance of the scalar row $z^{\top}\widetilde{h}_{t}$ from the span of $\{z^{\top}\widetilde{h}_{1},\dots,z^{\top}\widetilde{h}_{t-1}\}$, and $\widetilde{s}_{t}$ equals the squared-distance operator (in $\mathbb{R}^{m}$) from $\widetilde{h}_{t}$ to $\mathrm{span}\{\widetilde{h}_{1},\dots,\widetilde{h}_{t-1}\}$. Since $\widetilde{H}_{T}$ is block lower-triangular with diagonal block $US_{\mathcal{A}}$, and $US_{\mathcal{A}}$ has orthonormal rows (because $S_{\mathcal{A}}S_{\mathcal{A}}^{\top}=I_{m}$ and $U$ is orthogonal), the diagonal block is orthogonal to the past block-rows, implying
$\widetilde{s}_{t}\succeq I_{m}\quad\Rightarrow\quad\widetilde{s}_{t}^{-1}\preceq I_{m}.$ (H.14)
Combining (H.13)–(H.14) gives
$x^{\top}\widetilde{K}_{T}^{-1}x\leq\sum_{t=1}^{T}\|\widetilde{r}_{t}(x)\|_{2}^{2}.$ (H.15)
We now bound the residuals for $x=\widetilde{b}^{(R)}(c)$. As in the proof of Lemma G.2, the residual $\widetilde{r}_{t}(x)$ is the error of the best linear predictor of $x_{t}$ from the past $x_{1:t-1}$. Hence for any (possibly suboptimal) predictor $\widehat{x}_{t}$ measurable w.r.t. the past,
$\|\widetilde{r}_{t}(x)\|_{2}^{2}\leq\|x_{t}-\widehat{x}_{t}\|_{2}^{2}.$ (H.16)
Fix $t\geq 2$ and set the truncation length
$m_{t}:=\left\lceil\frac{\log t}{\log(1/\rho)}\right\rceil\quad\Rightarrow\quad\rho^{m_{t}}\leq\frac{1}{t}.$
We use the same predictor as in the proof of Lemma G.2, applied component-wise:
$\widehat{\widetilde{b}}^{(R)}_{t}:=\sum_{q=1}^{m_{t}}\beta_{q}\,\widetilde{b}^{(R)}_{t-q}(c),\qquad\beta_{q}:=(W^{q})_{ii},$ (H.17)
for an arbitrary but fixed $i\in\mathcal{A}$ (any $i$ works; only $|\beta_{q}|\leq 1$ is used).
Using same techniques as in the proof of Lemma G.2, i.e., using (i) $W^{q}R^{\ell-q}=R^{\ell}$, (ii) row-stochasticity, and (iii) $\|R^{k}\|_{2}\leq\rho^{k}$, gives the uniform tail bound
$\left\|\widetilde{b}^{(R)}_{t}(c)-\widehat{\widetilde{b}}^{(R)}_{t}\right\|_{2}\leq\frac{\sqrt{m}}{(1-\rho)^{2}}\,\frac{1}{t}.$ (H.18)
Here the extra factor $\sqrt{m}$ comes from bounding an $\ell_{2}$-norm over $m$ coordinates by $\sqrt{m}$ times a coordinate-wise bound.
Combining (H.16) and (H.18), we obtain
$\|\widetilde{r}_{t}(\widetilde{b}^{(R)}(c))\|_{2}^{2}\leq\left\|\widetilde{b}^{(R)}_{t}(c)-\widehat{\widetilde{b}}^{(R)}_{t}\right\|_{2}^{2}\leq\frac{m}{(1-\rho)^{4}}\,\frac{1}{t^{2}}.$ (H.19)
Using (H.15), summing over $t$ and using the bound (H.19) gives
$\left(\widetilde{b}^{(R)}(c)\right)^{\top}\widetilde{K}_{T}^{-1}\,\widetilde{b}^{(R)}(c)\leq\sum_{t=1}^{T}\|\widetilde{r}_{t}(\widetilde{b}^{(R)}(c))\|_{2}^{2}\leq\frac{m}{(1-\rho)^{4}}\sum_{t=1}^{\infty}\frac{1}{t^{2}}\leq\frac{\pi^{2}/6}{(1-\rho)^{4}}\,m.$
This proves the claim. ∎
Appendix I Extension to the Case When Observers’ Noise Does not Contribute (SecLDP)
We next extend the proof technique of Theorem 15 to the SecLDP setting, where the adversary observing a set $\mathcal{A}$ can remove the Gaussian noise injected by nodes in $\mathcal{A}$. Equivalently, one deletes the corresponding columns in each time-block column of the system matrix.
I.1 Proof of Theorem 16
We simply sketch the proof as it has exactly the same structure as the proof of Theorem 15.
In the SecLDP setting, the adversary observing $\mathcal{A}$ can remove the Gaussian noises injected by nodes in $\mathcal{A}$. This is modeled by right-multiplying each time-block column of the system matrix by
$P_{\mathcal{A}^{c}}:=I_{n}-\sum_{i\in\mathcal{A}}e_{i}e_{i}^{\top},\qquad\bar{H}_{T}:=H_{T}(I_{T}\otimes P_{\mathcal{A}^{c}}),\qquad\bar{K}_{T}:=\bar{H}_{T}\bar{H}_{T}^{\top}.$
For $j\notin\mathcal{A}$, $(I_{T}\otimes P_{\mathcal{A}^{c}})(c\otimes e_{j})=c\otimes e_{j}$, hence: $\bar{b}(c):=\bar{H}_{T}(c\otimes e_{j})=H_{T}(c\otimes e_{j})=b(c)$. Therefore the SecLDP sensitivity is
$\bar{\Delta}_{T}(j\to\mathcal{A})^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}\bar{K}_{T}^{+}\,b(c),$
and the difference to Theorem 15 is in the modified Gram matrix $\bar{K}_{T}$.
We decompose $W=P+R$ as in the single- and multi-node proofs, with $P=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$, $PR=RP=0$, and $\|R^{t}\|_{2}\leq\rho^{t}$. This gives a decomposition $\bar{H}_{T}=\bar{H}_{T}^{(P)}+\bar{H}_{T}^{(R)}$ and, expanding,
$\bar{K}_{T}=\bar{K}_{T}^{(PP)}+\bar{K}_{T}^{(RR)}+\bar{K}_{T}^{(PR)}+\bar{K}_{T}^{(RP)},\qquad\bar{K}_{T}^{(PP)}:=\bar{H}_{T}^{(P)}(\bar{H}_{T}^{(P)})^{\top},\ \ \bar{K}_{T}^{(RR)}:=\bar{H}_{T}^{(R)}(\bar{H}_{T}^{(R)})^{\top}.$
Unlike in the proof of Theorem 15, the cross terms $\bar{K}_{T}^{(PR)}+\bar{K}_{T}^{(RP)}$ do not vanish in general because of the right-multiplication by $P_{\mathcal{A}^{c}}$ (we cannot use the orthogonality $\mathbf{1}^{\top}R=0$). We keep these terms explicitly and control them by Cauchy–Schwarz: for any $x$ and any $\eta>0$,
$\big|x^{\top}(\bar{K}_{T}^{(PR)}+\bar{K}_{T}^{(RP)})x\big|\leq\eta\,x^{\top}\bar{K}_{T}^{(PP)}x+\eta^{-1}x^{\top}\bar{K}_{T}^{(RR)}x,$
so the cross terms can be absorbed into the stationary and transient terms.
The stationary term is computed explicitly. Using $P=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ and $P_{\mathcal{A}^{c}}\mathbf{1}=\mathbf{1}_{\mathcal{A}^{c}}$, one obtains the Toeplitz form
$\bar{K}_{T}^{(PP)}=\frac{n-m}{n^{2}}\,(J\otimes\mathbf{1}_{m}\mathbf{1}_{m}^{\top}),\qquad J_{t,s}=\min(t,s),$
i.e., compared to the proof of Theorem 15, we get the coefficient (nm)/n(n-m)/n because only the $\mathcal{A}^{c}$ noise coordinates remain. Consequently, the same backward-difference argument ($J^{-1}=D^{\top}D$) used in (H.11) gives the SecLDP linear term with slope $\frac{m}{n-m}$ (instead of $\frac{m}{n}$).
Finally, all terms involving $R$ (the pure transient contribution and the cross-term contribution) are controlled by the exactly the same Schur-complement technique as Lemma G.2 and Lemma H.1, giving a uniform bound independent of $T$:
$(b^{(R)}(c))^{\top}\bar{K}_{T}^{+}b^{(R)}(c)\leq C_{\mathrm{mix}}(W),$
and similarly for the cross terms. Putting the stationary and transient terms together gives the claim.
Appendix J Proofs for Section 5 (Utility Analysis)
J.1 From Absolute Spectral Gap to the Consensus Parameter $p$
We prove the standard contraction inequality and its implication for the parameter $p$ that is needed in the unified decentralized SGD analysis of (Koloskova et al., 2020).
Lemma J.1.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric and doubly-stochastic. Let $1=\lambda_{1}(W)\geq\lambda_{2}(W)\geq\cdots\geq\lambda_{n}(W)\geq-1$ be its eigenvalues. For any $X\in\mathbb{R}^{d\times n}$, define $\bar{X}:=X\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ and $\rho=\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}$. Then
$\|XW-\bar{X}\|_{F}^{2}\;\leq\;\rho^{2}\,\|X-\bar{X}\|_{F}^{2}.$ (J.1)
Consequently, the assumption
$\|XW-\bar{X}\|_{F}^{2}\leq(1-p)\|X-\bar{X}\|_{F}^{2}$ (J.2)
needed in the analysis of (Koloskova et al., 2020) holds with $p=1-\rho^{2}$.
Proof.
Let $P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ be the orthogonal projector onto $\mathrm{span}\{\mathbf{1}\}$. Then $\bar{X}=XP$ and, since $W\mathbf{1}=\mathbf{1}$ and $\mathbf{1}^{\top}W=\mathbf{1}^{\top}$, we have $WP=PW=P$. Hence
$XW-\bar{X}=XW-XP=X(W-P)=(X-\bar{X})(W-P),$
because $\bar{X}(W-P)=XP(W-P)=X(PW-P)=0$. Now, WPW-P is symmetric with eigenvalues 0 (along $\mathbf{1}$) and $\lambda_{2}(W),\ldots,\lambda_{n}(W)$ on $\mathbf{1}^{\perp}$, so $\|W-P\|_{2}=\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}=\rho$ (since $W$ is symmetric). Thus
$\|XW-\bar{X}\|_{F}=\|(X-\bar{X})(W-P)\|_{F}\leq\|X-\bar{X}\|_{F}\,\|W-P\|_{2}=\rho\,\|X-\bar{X}\|_{F}.$
Using the relation $1-\rho^{2}=(1+\rho)(1-\rho)$ we directly get the following auxiliary result.
Lemma J.2.
Let $W$ be symmetric and doubly-stochastic and define the absolute spectral gap
$\gamma:=1-\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}\in(0,1].$
With $p=1-\rho^{2}$,
$\frac{1}{p}=O\!\left(\frac{1}{\gamma}\right),\;\;\frac{1}{p^{2}}=O\!\left(\frac{1}{\gamma^{2}}\right).$ (J.3)
J.2 Deriving the Utility Bound from Results of (Koloskova et al., 2020)
We derive Theorem 17 by plugging our privacy calibration into a strongly-convex convergence bound by (Koloskova et al., 2020). More specifically, we use the following template bound (a specialization of the unified analysis to $\tau=1$ and strongly convex objectives).
Lemma J.3 (Koloskova et al. 2020).
Assume each $f_{i}$ is $L$-smooth and $\mu$-strongly convex. Let $\tau=1$ (full-batch gradients). Assume unbiased stochastic gradients with variance proxy $\bar{\sigma}^{2}$ and heterogeneity $\bar{\zeta}^{2}$. Assume the consensus condition (J.2) holds with parameter $p\in(0,1]$. Then, for a constant step size $\eta\simeq 1/L$, the averaged iterate $\bar{\theta}^{t}$ satisfies
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{\theta}^{t})-f^{\star}\right]\;\leq\;\widetilde{O}\!\left(\frac{\bar{\sigma}^{2}}{n\mu T}\;+\;\frac{L(\bar{\zeta}^{2}+p\bar{\sigma}^{2})(1-p)}{\mu^{2}p^{2}T^{2}}\;+\;\frac{L\|\bar{\theta}^{0}-\theta^{\star}\|_{2}^{2}}{p}\exp\!\left(-c_{0}\,\frac{\mu p}{L}T\right)\right),$ (J.4)
[…]
J.3 Proof of Theorem 17
Proof.
To specialize Lemma J.3 to our setting, we start by identifying the gradient noise as the DP noise. Namely, under (5.1) with full gradients and $\xi_{i}^{t}\sim\mathcal{N}(0,\sigma_{\mathrm{DP}}^{2}I_{d})$, the only stochasticity is the DP noise, and hence
$\sigma_{i}^{2}:=\mathbb{E}\|\xi_{i}^{t}\|_{2}^{2}=d\,\sigma_{\mathrm{DP}}^{2},\qquad\bar{\sigma}^{2}=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}^{2}=d\,\sigma_{\mathrm{DP}}^{2}.$
Next, we adjust Gaussian noise to $(\varepsilon,\delta)$-DP guarantees by (see, e.g., Dwork and Roth, 2014)
$\sigma_{\mathrm{DP}}\;=\;\frac{\Delta_{T}(j\to i)\sqrt{2\log(1.25/\delta)}}{\varepsilon}.$
As shown in the main text, using Thm. 15, we have
$(\Delta_{T}(j\to i))^{2}\;\leq\;\frac{T}{n}+\frac{2}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{2}{\gamma^{4}}$ (J.5)
and by adjusting the Gaussian noise via the standard condition (Dwork and Roth, 2014)
$\sigma_{\mathrm{DP}}\;=\;\frac{\Delta_{T}(j\to i)\,\sqrt{2\log(1.25/\delta)}}{\varepsilon}$ (J.6)
we get
$\sigma_{\mathrm{DP}}^{2}\;\lesssim\;\frac{\log(1/\delta)}{\varepsilon^{2}}\left(\frac{T}{n}+\frac{1}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{1}{\gamma^{4}}\right).$ (J.7)
For symmetric doubly-stochastic $W$, Lemma J.1 gives $p=1-\gamma^{2}$ and Lemma J.2 implies $p\geq\gamma$ and $1/p=O(1/\gamma)$, $1/p^{2}=O(1/\gamma^{2})$.
We now bound each term in (J.4).
(i) Leading term. Using (J.7),
$\frac{\bar{\sigma}^{2}}{n\mu T}$ $=\widetilde{O}\!\left(\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n^{2}\,\mu}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{2}\sqrt{T}}\;+\;\frac{d\,\log(1/\delta)}{\varepsilon^{2}\,n\,\mu}\cdot\frac{1}{\gamma^{4}T}\right)$ (J.8)
(ii) Heterogeneity term. We split:
$\frac{L(\bar{\zeta}^{2}+p\bar{\sigma}^{2})(1-p)}{\mu^{2}p^{2}T^{2}}\;\leq\;\frac{L\bar{\zeta}^{2}}{\mu^{2}p^{2}T^{2}}\;+\;\frac{L\bar{\sigma}^{2}}{\mu^{2}pT^{2}},$
since $(1-p)\leq 1$ and $(p\bar{\sigma}^{2})(1-p)/p^{2}\leq\bar{\sigma}^{2}/p$. Using $p\geq\gamma$, we have that
$\frac{L\bar{\zeta}^{2}}{\mu^{2}p^{2}T^{2}}$ $\leq\frac{L\bar{\zeta}^{2}}{\mu^{2}\gamma^{2}T^{2}}$ (J.9)
and
$\frac{L\bar{\sigma}^{2}}{\mu^{2}pT^{2}}$ $=\widetilde{O}\!\left(\frac{L}{\mu^{2}\gamma T^{2}}\cdot\frac{d\log(1/\delta)}{\varepsilon^{2}}\left(\frac{T}{n}+\frac{1}{\gamma^{2}}\sqrt{\frac{T}{n}}+\frac{1}{\gamma^{4}}\right)\right)$
$=\widetilde{O}\!\left(\frac{dL\log(1/\delta)}{\varepsilon^{2}\mu^{2}}\cdot\frac{1}{n\gamma T}\right).$ (J.10)
(iii) Transient term. Using $p\geq\gamma$,
$\frac{L\|\bar{\theta}^{0}-\theta^{\star}\|_{2}^{2}}{p}\exp\!\left(-c_{0}\frac{\mu p}{L}T\right)\;\leq\;\frac{L\|\bar{\theta}^{0}-\theta^{\star}\|_{2}^{2}}{\gamma}\exp\!\left(-c\,\gamma\frac{\mu}{L}T\right)$ (J.11)
for some constant $c>0$.
Combining (J.8), (J.9) and (J.11) gives the statement of Theorem 17. ∎